### v1

# Mitigating Heterogeneity in Federated Multimodal Learning with Biomedical Vision-Language Pre-training

Abstract
abstract Vision-language pre-training (VLP) has arised as an efficient scheme for multimodal representation learning, but it requires large-scale multimodal data for pre-training, making it an obstacle especially for biomedical applications. To overcome the data limitation, federated learning (FL) can be a promising strategy to scale up the dataset for biomedical VLP while protecting data privacy. However, client data are often heterogeneous in real-world scenarios, and we observe that local training on heterogeneous client data would distort the multimodal representation learning and lead to biased cross-modal alignment. To address this challenge, we propose Federated distributional Robust Guidance-Based (FedRGB) learning framework for federated VLP with robustness to data heterogeneity. Specifically, we utilize a guidance-based local training scheme to reduce feature distortions, and employ a distribution-based min-max optimization to learn unbiased cross-modal alignment. The experiments on real-world datasets show our method successfully promotes efficient federated multimodal learning for biomedical VLP with data heterogeneity.
Vision-Language Pre-training, Federated Learning, Data Heterogeneity
abstract Vision-language pre-training (VLP) has arised as an efficient scheme for multimodal representation learning, but it requires large-scale multimodal data for pre-training, making it an obstacle especially for biomedical applications. To overcome the data limitation, federated learning (FL) can be a promising strategy to scale up the dataset for biomedical VLP while protecting data privacy. However, client data are often heterogeneous in real-world scenarios, and we observe that local training on heterogeneous client data would distort the multimodal representation learning and lead to biased cross-modal alignment. To address this challenge, we propose Federated distributional Robust Guidance-Based (FedRGB) learning framework for federated VLP with robustness to data heterogeneity. Specifically, we utilize a guidance-based local training scheme to reduce feature distortions, and employ a distribution-based min-max optimization to learn unbiased cross-modal alignment. The experiments on real-world datasets show our method successfully promotes efficient federated multimodal learning for biomedical VLP with data heterogeneity.
Figure 1: Federated multi-modal pre-training aims to de-centralizedly utilizing private dataset to tackle the data hungry problem. Fitting heterogeneous local datasets that correspond to different distributions of latent semantics (e.g. disease) would result in biased models. Our FedRGB framework has successfully upgraded multi-modal pretraining methods by introducing robust alignment module as guidance during local training time.
1 Introduction
Vision-language pre-training (VLP) learns efficient multimodal representations by extracting latent semantics from a large amount of image-text pairs, where the dataset scale largely impacts the performance of the learned model (Oquab et al., 2023). However, enlarging pre-training dataset is a non-trivial challenge especially for biomedical applications which is restricted by privacy issues of clinical and patient data (Ladbury et al., 2023). To utilize data across different institutes, recent work (Lu et al., 2023a) has attempted to introduce federated learning into the biomedical VLP domain. However, in real-world scenarios, datasets collected from different institutes are always heterogeneous, which is a long-standing practical challenge in federated learning (Ghosh et al., 2019; Zhang et al., 2023a; Huang et al., 2022). For example, one hospital receives a high proportion of patients with pneumonia, while another hospital may undertake more patients with tuberculosis, due to variety of patients cohort and geographics. While the data heterogeneity would essentially damage the performance of the federally trained model, current FL methods for biomedical VLP (Lu et al., 2023a) haven’t taken this vital factor into consideration to provide an effective solution. To address this issue, simply averaging client models trained on heterogeneous local datasets results in biases as shown in Figure 1. Firstly, the client model would be heterogeneous, and re-training on the sent-back server model might distort the aggregated structure of the representation space. Secondly, averaging model parameters cannot produce a robust alignment between the two modalities, which not only affects the generalization ability, but also prevents the model from learning more diverse correlations between vision and language. Furthermore, as we will demonstrate in our empirical findings (Sec 5.2), heterogeneity harms the performance of pre-trained model by affecting the deep layers which serve as an alignment module in the VLP model. In this paper, we propose the FedRGB framework to mitigate heterogeneity in federated biomedical VLP to enable its deployment in real-world scenarios. To alleviate the distortion in local client training, we introduce a teacher alignment module to provide unbiased cross-modal alignment by generating anchor representation space for regularization. We employ a distributionally robust optimization (DRO)-based algorithm to learn a robust teacher alignment module, which is optimized to better alleviate the worst-case distortion on the feature encoder caused by local cross-modal alignment. Our contributions primarily focus on:
  • We for the first time tackle the problem of biomedical VLP under the federated setting by utilizing heterogeneous datasets from different institutes. We conduct empirical studies to analyze the influence of data heterogeneity on federated multimodal learning.
  • We propose FedRGB framework to mitigate the heterogeneity in federated multimodal learning. Specifically, our method introduces the guidance-based local training scheme for learning unbiased cross-modal alignment to alleviate representaion distortions.
  • Experiment results show the effectiveness of our method in learning multimodal representations under the federated setting by appling to various downstream tasks including image-text retrieval, classification, and segmentation.
2 Related Work
2.1 Vision-Language Pre-training
Vision-Language Pre-training (VLP): Leveraging data of different modalities to enable multimodal reasoning and representation learning for various cross-modal tasks has become an active research field in machine learning (Radford et al., 2021). Pre-training vision-language models on large datasets and then transferring learned knowledge to downstream tasks has become a popular approach to extract semantics from unlabeled image-text paired datasets (Li et al., 2022a; Bao et al., 2022). To find the shared latent space, current works often focus on learning generalizable and robust alignment relationships between the two modalities (Radford et al., 2021; Chen et al., 2020a), which measures the ability to connect and retrieve the representation of the paired data. Biomedical Vision-Language Pre-training: Most of the existing works in biomedical vision-language pre-training are based on the twin modality-specific encoders for medical images and reports respectively, with the image-report contrastive loss (Zhang et al., 2022b) to align visual and language representations on a joint latent space. A significant improvement has been achieved using the local contrastive loss which aims to align fine-grained representations (Huang et al., 2021). Incorporating medical domain knowledge, Wang et al. (2022) proposes a well-designed method that aligns the image and report embeddings in three different granularities. Recent works have also tried to address other practical issues of misalignments (Bannur et al., 2023). However, most of these methods haven’t considered the adaptation to de-centralized scenarios during pre-training and thus might fail to handle the potential damage from data heterogeneity.
2.2 Federated Learning
Heterogeneous Federated Learning: Federated learning aims to de-centralize model training by aggregating models trained on local clients to obtain more generalizable models with scaled-up training datasets while protecting data privacy (McMahan et al., 2017), which achieved success in many applications especially those with data privacy issues like biomedical data. However, client datasets are heterogeneous in most cases, which damages the performance of the aggregated model (Ghosh et al., 2019). Existing methods have tried to solve the data heterogeneity problem in many ways (Li & Wang, 2019; Collins et al., 2021). However, most methods either require additional communications on representations (Zhang et al., 2023a) which are not privacy-preserved and inefficient, or need large-scale public datasets for server aggregation (Huang et al., 2022; Yu et al., 2023b) that does not fit the case for biomedical data. Besides, remaining methods are often designed for supervised tasks that focus on label distribution (Diao et al., 2020) or feature distribution (Li et al., 2022b), which cannot be directly adapted to the vision-language pre-training task. While the self-supervised pre-trained models can be biased due to the heterogeneous datasets in the biomedical domain (Krishnan et al., 2022) and over-reliance on spurious features (Saab et al., 2022), distributionally robust optimization (Deng et al., 2020) can optimize the worst-case performance thus alleviating these issues (Liu et al., 2022). Recent works have considered its application in unsupervised and self-supervised learning senarios (Han et al., 2023; Zhang et al., 2022a; Capitani et al., 2024). Federated Multi-modal Learning: Federated learning strategies have been extended to multimodal datasets in recent works (Zhou & Wang, 2022), with the goal to reduce training cost and promote personalized deployment of multimodal foundation models (Su et al., 2022; Chen et al., 2023b; Lu et al., 2023b). For example, Yu et al. (2023a) considers a client-server collaborative pre-training setting, benefiting the pre-training process of the server model by leveraging representations from local modalities. Nevertheless, these works require a pre-training stage in the server while ignoring that only on client datasets with limited samples. The closest work to ours is Lu et al. (2023a), which adopts federated learning to scale up the pre-training dataset. However, this work lacks generalization ability and hasn’t considered the data heterogeneity issue that is very common and significant in practice.
3 Problem Formulation
We aim at mitigating the data heterogeneity issues in federally pre-training biomedical vision-language models to better capture the semantic alignment between the two modalities, with heterogeneous private datasets from different clients. In this section, we will formulate our problem set-up and assumptions. Please see Appendix A for the detailed preliminary. Federated Vision-Language Pre-training. In this paper, we consider the federated vision-language pre-training (VLP) problem in the biomedical domain. Suppose we are given $N$ clients $\{C_{i}\}_{i=1}^{N}$, each client has a corresponding local yet private dataset $\{D_{i}\}_{i=1}^{N}$ that can’t be easily shared due to privacy issues. Without access to centralized pre-training, federated pre-training aims to conduct vision-language representation learning across local clients, while still capture a robust and generalizable cross-modal alignment, by efficiently aggregating client models $\{M_{i}\}_{i=1}^{N}$ trained on their private datasets. Similar to conventional federated learning scenarios, our federated pre-training consists of multiple communication turns. For communication turn $t$, client models $\{M_{i}\}_{i=1}^{N}$ completed the same number of local iteration steps respectively, and the server receives and aggregates the weights sent from local models to produce the aggregated server model $M^{t}$. Then it will be sent back to clients and re-trained on $\{D_{i}\}_{i=1}^{N}$ separately to get new local models $\{M_{i}^{t+1}\}_{i=1}^{N}$. The federated vision-language pre-training is a more practical and general task compared to related works. Table 1 lists the most similar settings in previous works, and the difference between our task and theirs.
Table 1: Comparison between related multi-modal federated learning tasks and our federated biomedical vision-language pre-training task. We organize the related works according to our key areas of interest. (1). Heterogeneous clients. (2). Multi-modal clients. (3). Transferability. (4). Prior model-free. (5). No communication on public data or feature representations.
Task hetero. m.m.c. transf. no prior no feat. com.
(Yu et al., 2023a) $\checkmark$ $\checkmark$
(Lu et al., 2023b) $\checkmark$ $\checkmark$ $\checkmark$
(Chen et al., 2023b) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
(Lu et al., 2023a) $\checkmark$ $\checkmark$ $\checkmark$
FedRGB $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Data Formulation. We consider the multimodal datasets including vision and language modalities $X,Y$, consisting of respective data samples $X=\{x_{i}\}_{i=1}^{n},Y=\{y_{i}\}_{i=1}^{m},n,m\in N^{+}$. We assume there exists latent variables $U$ that embed semantics across two modalities. For example, the disease category of a given X-ray radiology-report pair is a latent variable connecting these two modalities, which determines the pathology region of the radiology image and semantics in the diagnosis report. For heterogeneous local datasets, we assume the distribution of latent variables $P_{i}(U)$ is different across clients.
4 Method
Figure 2: (a) Our proposed FedRGB upgrades backbone multi-modal pre-training strategies via introducing a robust guidance to local pre-training. This approach can reduce the distortion on feature encoders and learn better cross-modal alignment. (b) We employ a distribution-based min-max optimization to learn robust cross-modal alignment, which will be used as guidance in local pre-training.
We propose a de-centralized framework to mitigate the data heterogeneity in federated biomedical vision-language pre-training, for learning an efficient vision-language model that can robustly capture reliable cross-modal semantic alignments. In this section, we first introduce our approach with a guidance-based local training strategy to reduce the distortion on encoders caused by fitting client datasets. Secondly, we will illustrate how our approach employ a distribution-based min-max optimization to learn unbiased cross-modal alignment. Algorithm 1 shows the pipeline of our method.
4.1 Guidance-based Local Client Training
As shown in our empirical findings (Sec. 5.2), biomedical vision-language models pre-trained on heterogeneous client datasets are affected by the local correlations of the two modalities, which can’t be addressed by simply averaging the weights of local models. Furthermore, this mainly falls on the deep layers which act as alignment modules, where the heterogeneous alignment module would distort the learned encoder during local iterations as shown in Sec. 7. Inspired by these findings, we introduce a teacher aligner to guide the representation space learned by local models to converge to a generalizable and robust one. Teacher alignment module. To alleviate the distortion and bias when learning local features encoders and alignment modules, we propose to utilize a teacher alignment module to generate anchor representations conditioned on the output feature vector of the encoder to encourage local models to learn a homogeneous representation space across clients. Suppose after one client-server communication turn, given the local dataset of client $c\in\mathcal{C}$ sampled from $(x^{c},y^{c})\sim D_{c}$, and teacher alignment module $g_{v}^{*},g_{l}^{*}$ corresponding to vision and language modality, we target at updating the feature encoders of image and text $f_{v},f_{l}$ and the corresponding alignment modules $g_{img},g_{txt}$, by training on the local client dataset $D_{c}$, under the guidance of teacher alignment modules. Guiding local training with robust aligner. As shown in Fig. 2, given the image-text pairs from local dataset as input $(x^{c},y^{c})\sim D_{c}$, we denote the feature vectors output from the feature encoder modules as $(v^{c},l^{c})$. Then getting through the alignment module, the aligned representations output from local aligners are denoted by $(z_{I}^{c},z_{T}^{c})$, while those from teacher aligners are $(z^{c*}_{I},z^{c*}_{T})$. The goal is to reduce the discrepancy between two sets of aligned representations produced by local and teacher aligners, so that we can keep locally trained model will not be misled by the bias from local datasets and all local clients can converge to learn a similar embedding space that captures globally reliable and generalizable cross-domain alignment. The teacher guidance loss can be formulated as below:
$\footnotesize\mathcal{L}_{guide}=\sum_{i=1}^{N_{c}}(z^{c*}_{I,i}-z^{c}_{I,i})^{2}+\sum_{i=1}^{N_{c}}(z^{c*}_{T,i}-z^{c}_{T,i})^{2}$ (1)
, where $N_{c}$ refers to the number of samples in client the $c$. We combine the guidance loss with the optimization goal $\mathcal{L}_{task}$ of the initial biomedical vision-language pre-training method, which includes e.g. image-text contrastive loss in ConVIRT (Zhang et al., 2022b). The overall loss function for local client training can be written as:
$\footnotesize\mathcal{L}_{local}=\mathcal{L}_{guide}+\alpha\mathcal{L}_{task}$ (2)
where $\alpha$ is the hyper-parameter that balances between guidance regularization and training task in local client updating.
4.2 Learning Unbiased Cross-modal Alignment via Distribution-based Optimization
In order to guide the local client training, we need to provide a teacher alignment module as mentioned in the last section, which can capture globally unbiased and generalizable cross-domain alignment. How can we obtain such a teacher guidance model? We propose to learn the unbiased cross-domain alignment by considering the robustness across various distributions of local clients. Detailed problem formulation and description could be seen in Appendix B. A teacher alignment module should avoid distorting the aggregated feature encoder while learning the unbiased cross-modal alignments. Additionally, we introduce a much harder task to force the module to be more robust to biased alignment, by considering the influence of the biased alignment module on the feature encoder. Furthermore, to learn a teacher model that performs well on varies data distributions, we minimize distortion degree under the most biased local alignment with the “worst-case” local data distribution. Learning robust cross-modal alignment through distributionally robust optimization. Inspired by the distributionally robust optimization (DRO) (Cotter et al., 2019), we model the de-centralized pre-training framework as a two-stage DRO learning process. In the first stage, the local models are trained with $\mathcal{L}_{local}$ (Eq. 2), while the whole model is jointly optimized with the guidance of the teacher alignment module. In the second stage, we employ the decentralized DRO approach to learn a robust teacher alignment module that performs well conditioned on the worst-case local models where the feature encoder is distorted the most severely. Learning object of alleviating distortion in local triaing. After finishing the local training in each communication round, we aim to employ an extra de-centralized learning process in the client-side to optimize the alignment module on the worst-case distribution of an uncertainty set constructed by the local aligner and local data distribution, denoted as $P_{i}$ for simplicity. The uncertainty set can be written as:
$\footnotesize Q^{c}:\{D(Q||P^{c})\leq\rho\},P^{c}=\sum_{i=1}^{N}w_{i}P_{i},w\in\Delta_{N-1}$ (3)
where $N$ is the number of clients, $\Delta$ is a $(N-1)$-dim simplex, $D$ is a discrepancy measure, and $\rho$ is the uncertainty radius that measures the uncertainty degree of the unseen distributions and influences model’s generalization ability. The details of the $P_{i}$ can be seen in Appendix B. With the teacher alignment module denoted as $g^{*}$ and the aggregated encoder $f^{*}$ after the first stage training, the optimization goal can be constructed by distributions $P_{i}$ of the support set:
$\footnotesize min_{g^{*}}\ max_{\textbf{w}\in\Delta_{N-1}}\sum_{i=1}^{N}w_{i}\ \mathcal{L}_{dro}(g^{*},P_{i},f^{*}),s.t.D(N\textbf{w}||\textbf{1})\leq\rho$ (4)
,where $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$ are the losses calculated on the local clients respectively. While we learn $g^{*}$ by gradient descent, employing an alternative optimization approach, the mirror gradient descent process of optimizing w could be solved efficiently (Zhang et al., 2023b; Duchi et al., 2008):
$\footnotesize{w_{i}^{(r+1)}}=\text{Proj}_{\mathcal{Q}{c}}\left({w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}}/\sum_{i=1}^{N}{w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}},\rho\right)$ (5)
where $\text{Proj}_{\mathcal{Q}{c}}$ refers to the projection to $\{w:D(N\textbf{w}||\textbf{1})\leq\rho\}$, $v_{i}$ represents the loss $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$, $r$ refers to the communication turn.
Finally, we formulate $\mathcal{L}_{dro}$ to measure the performance of the teacher alignment module of alleviating distortion on the encoder. Based on Eqs 3 and 4, we update the teacher alignment module in local clients with weights w. Specifically, given a client $c$, we use $f^{*}$ learned in the first stage of the current communication turn to generate features $(\tilde{v^{c}},\tilde{l^{c}})$ which are not distorted by the local aligner during the second stage’s training. And a learnable copy of $f^{*}$ to generate distorted features $(v^{c},l^{c})$. The distorted features are input to the frozen local aligner $g$ to calculate the task loss $L_{task}$, and are also input to the teacher aligner $g^{*}$ to get calibrated representations $(z_{I}^{c*},z_{T}^{c*})$. Meanwhile, $(\tilde{v^{c}},\tilde{l^{c}})$ are input to $g^{*}$ to get representation $\tilde{(z_{I}^{c*}},\tilde{z_{T}^{c*}})$ as anchors for $(z_{I}^{c*},z_{T}^{c*})$ projected by teacher aligner using distorted features, for regularization. The learning object of the second stage could be written as:
$\footnotesize\mathcal{L}_{dro}=\beta(\sum_{i=1}^{N_{c}}(\tilde{z_{I,i}^{c*}}-z_{I,i}^{c*})^{2}+\sum_{i=1}^{N_{c}}(\tilde{z_{T,i}^{c*}}-z_{T,i}^{c*})^{2})+\mathcal{L}_{CL}+\mathcal{L}_{task}$ (6)
,where $\mathcal{L}_{CL}$ is the contrastive loss (Radford et al., 2021) between $\tilde{(z_{I}^{c*}},\tilde{z_{T}^{c*}})$, $\mathcal{L}_{task}$ refers to the loss function of the backbone pre-training method and $N_{c}$ refers to the number of samples in the client $c$.
5 Experiment
5.1 Experiment Setting
5.1.1 Pre-training setup
Following (Wang et al., 2022), we utilize the MIMIC-CXR (Bigolin Lanfredi et al., 2022) dataset for biomedical vision-language pre-training. During the de-centralized pre-training process, local clients only have access to their private datasets. Following Yan et al. (2023), we employ the Latent Dirichlet Allocation (LDA) (Blei et al., 2003) to divide the MIMIC-CXR dataset based on a 5-way disease attribute to construct the heterogeneous client datasets. We set the heterogeneity degree in the LDA algorithm to be 1. The client datasets are highly heterogeneous as shown in Figure 3 and the appendix. We divide the MIMIC-CXR into 5 heterogeneous subgroups to construct 5 clients. Each divided dataset consists of train splits and validation splits based on the notation of the MIMIC-CXR. For the pre-training reported in the main table, we set the number of communications to 25, and we randomly sampled 50 batches of data from the client datasets for each communication. For more details, please see the appendix. For analysis experiments on out-of-distribution generalization ability, we utilize the EHRXQA (Bae et al., 2023) to evaluate the transferability of the pre-trained model to the unseen domain.
Figure 3: The descriptive statistic of the divided 5 subgroups, it can be seen that they are highly heterogeneous.
5.1.2 Downstream tasks
We conduct several downstream tasks following (Wang et al., 2022) to evaluate the generalization ability of the pre-trained model. For more details, please see the appendix. Few-shot classification.We test their performance on multiple image classification benchmarks: (1) RSNA Pneumonia Detection (RSNA) (Shih et al., 2019). (2) Covidx (Wang et al., 2020). We fine-tune our pre-trained model with an additional linear layer on 1%, 10% percentage of the training dataset, and evaluate the classification accuracy. Medical image segmentation. We conduct experiments following (Wang et al., 2022) on medical image segmentation on the RSNA (Wang et al., 2020) benchmark. We freeze the encoder and fine-tune a U-Net decoder using 1%, 10% of the training data, and then use the Dice score for evaluation. Image retrieval. To verify the pre-trained models have learned semantics and cross-modal alignments, we conduct the image retrieval task. Specifically, we utilize the validation splits of the clients for evaluation, these datasets are unseen in pre-training. We utilize the top-1 recall accuracy and top-5 recall accuracy to evaluate the pre-trained models.
5.1.3 Backbones and Baselines
We focus on upgrading biomedical vision-language pre-training (VLP) strategies to heterogeneous federated learning scenes. Since federated heterogeneous VLP is a new topic, we adopt existing biomedical VLP baselines to federated strategies (Lu et al., 2023a). Specifically, we consider adopting simple language-image contrastive alignment (ConVIRT) (Zhang et al., 2022b; Radford et al., 2021), global-local language-image contrastive alignment(GLoRIA) (Huang et al., 2021), and Multi-Granularity Cross-modal Alignment (MGCA) (Wang et al., 2022). Since the federated multi-modal per-training is a novel setting, we adapted FedMAE (Yan et al., 2023), FedEMA (Zhuang et al., 2022), and FedX (Han et al., 2022) for comparison. These are self-supervised learning methods in the federated learning domain which also focus on tackling the data heterogeneity. For basic federated learning baselines, we consider simple averaging (FedAvg) (McMahan et al., 2017; Lu et al., 2023a), centralized training (Local), and decentralized training (Global). For fair comparisons, we combine each federated learning strategy with different biomedical VLP methods. For baselines pre-trained in Local strategy, we report the averaged performance of the local models. For fair comparisons, we re-implement the methods with the same backbones. To enable uni-modal self-supervised learning baselines to be applied to our setting, we employ the image-text contrastive loss with the same hyperparameters to our method for fairness. We utilize a ViT-base (Dosovitskiy et al., 2020) as our vision encoder and utilize the Bert-base (Devlin et al., 2018) as our text encoder. We pre-process the input following (Wang et al., 2022). We utilize an additional transformer layer from ViT-base and Bert-base as the align module for vision and language respectively.
5.2 Empirical Finding
In this section, we will demonstrate our key empirical findings for federated multi-modal pre-training under heterogeneous client datasets, which actually motivates us to propose our Federated distributional Robust Guidance-Based (FedRGB). We conduct experiments on the image-text retrieval task, which reflects the ability to capture vision-language alignments. We utilize the ConVIRT backbone strategy and consider the performance of the centralized pre-trained model (Global), the server model pre-trained in FedAvg(FedAvg), and the average performance of decentralized pre-trained local models (Local.avg.).
Figure 4: The comparison of retrieval acc. on each client denoted as $\{C_{i}\}_{i=1}^{5}$,of centralized, FedAvg, and averaged acc. of decentralized pre-trained models.
Figure 5: The performance of the server model after 25 com. turns and the averaged performance of corresponding client models after 25 and 26 com. turns, on each client denoted as $\{Client_{i}\}_{i=1}^{5}$.
Figure 6: The averaged acc. on the test set of each client. We show the performance of FedAvg pre-trained baseline and its retrained models on different client datasets shown as $\{C_{i}\}_{i=1}^{5}$.
Figure 7: The averaged acc. on each client. We show the acc. of centralized and FedAvg pre-trained baselines, and de-centralized pre-trained models shown as $\{C_{i}\}_{i=1}^{5}$ retrained on the union of training splits of client datasets.
FedAvg can aggregate semantics from local clients, but the performance would be damaged by data heterogeneity. Figure 4 shows retrieval accuracies of the strategies we have considered, where the FedAvg pre-trained model achieves higher accuracies than those of decentralized ones with a large margin. However, the centralized pre-trained model acts as an upper bound and there still exists a great space for improvement for FedAvg. Table 12 in the appendix shows the detailed result. Further, we examine if updating the aggregated server model in FedAvg would result in distortion in the learned representation space. Retraining the server model on local datasets would distort its feature encoders. Figure 5 demonstrates the effect of distortion by fitting local cross-modal alignments of the corresponding client datasets. We consider a set of models that finished the 25th local iteration and their corresponding aggregated server model. While the average accuracy of local models is less than their aggregated model, the average performance of models re-trained on local clients starting from the aggregated model degrades a lot. Table 13 in the appendix shows the detailed results. Diverse information is aggregated in shallow layers that mainly extract feature semantics. We re-train the first 4 shallow layers of the aggregated server model on local clients’ datasets for 500 iteration steps. As shown in figure 6, the performance of retrained models decreases a lot, which means shallow layers that extract feature semantics are distorted by heterogeneous local datasets. Table 17 in the appendix shows the detailed results. De-centralized pre-trained deep layers that act as the align module is harmful for extracting diverse semantics. We re-train the first 4 shallow layers of the decentralized pre-trained model on the union of client datasets for 500 steps. Although the performance has successfully increased, there is still a large gap between that of the FedAvg pre-trained baselines and them as shown in Figure 7. Therefore, the frozen deep layers prevent the model from learning more diverse semantics from the scaled dataset. Table 18 in the appendix shows the detailed results. Overall, from the empirical findings, we conclude that federated multimodal pre-training is sensitive to the data heterogeneity which results in learning distorted cross-modal alignment. Simply averaging the local model weights does not solve this issue essentially. Thus, we are motivated to propose our FedRGB.
Table 2: Downstream task performance. We report the few-shot classification accuracy on Covid and RSNA, the retrieval accuracy on Chest imaGenome, and the Dice score on RSNA.
Strategy RSNA (cls.) Covid (cls.) RSNA (seg.) Chest imaG. (ret.)
$1\%$ $10\%$ $1\%$ $10\%$ $1\%$ $10\%$ Rec.@1 Rec.@5 Wst.@1 Wst.@5
Decentralized 82.27 82.83 79.45 88.25 67.44 70.75 15.51 51.12 13.62 45.97
FedX        (Han et al., 2022) 80.16 81.17 80.50 89.25 71.27 71.61 3.22 16.76 1.58 11.94
FedMAE (Yan et al., 2023) 76.97 80.66 42.25 49.00 69.48 73.38 5.05 23.91 4.34 20.86
FedEMA (Zhuang et al., 2022) 82.81 83.19 79.25 86.50 70.94 73.65 24.01 67.05 21.85 62.44
FedAvg   (Lu et al., 2023a) 83.19 83.31 78.00 88.50 69.59 71.50 28.83 72.12 25.34 66.68
FedRGB (Ours) 83.21 83.54 82.50 90.00 71.51 74.08 29.49 73.03 26.99 67.65
Centralized 83.39 84.61 82.50 92.00 72.60 76.35 41.45 84.22 38.55 79.98
Table 3: Downstream task performance on different multi-modal pre-training backbone methods.
Strategy Method RSNA (cls.) Covid (cls.) RSNA (seg.) Chest imaG. (ret.)
$1\%$ $10\%$ $1\%$ $10\%$ $1\%$ $10\%$ Rec.@1 Rec.@5 Wst.@1 Wst.@5
Decentralized GLoRIA 82.34 82.89 77.90 88.75 71.07 72.13 17.15 52.50 15.19 48.73
FedAvg GLoRIA 83.16 83.28 77.50 89.00 71.37 72.38 29.90 73.80 27.78 69.46
FedRGB (Ours) GLoRIA 83.39 83.51 78.50 89.25 72.43 73.50 30.94 73.89 27.86 69.53
Centralized GLoRIA 84.01 84.74 82.25 91.75 73.63 73.64 41.69 84.04 39.00 80.52
Decentralized MGCA 81.90 82.71 77.80 87.55 62.77 70.20 15.21 50.44 13.37 45.39
FedAvg MGCA 82.63 83.54 75.75 88.25 70.06 71.39 29.33 73.71 26.81 70.49
FedRGB (Ours) MGCA 82.76 83.61 77.75 88.25 70.71 72.52 29.96 74.15 27.11 69.53
Centralized MGCA 83.96 84.49 79.50 89.50 70.71 72.52 39.89 83.50 36.88 80.28
5.3 Main Result
Our method learns robust and enriched cross-modal alignment and has better transferability. Table 2 shows the image-text retrieval results. Here we utilize the ConVIRT as the backbone pre-training method. We report the average recall accuracy and the worst-case recall accuracy. The two sets of acc. are higher than those of FedAvg, which means our model can not only capture more cross-modal alignment but optimize the performance on the worst client dataset, which verifies the robustness and fairness of our DRO-based strategy. Table 2 shows the results of the classification and segmentation downstream tasks. In general, our method has better transferability than the traditional FedAvg strategy, most of the compared metrics of our method are higher than those of the baseline, and are closer to those of the centralized pre-trained model, which is the upper bound baseline. Other self-supervised baselines that aim to tackle the data heterogeneity have weaker performance compared to FedAvg, we analyze that this is due to the lack of the modeling of heterogeneous multi-modal alignment. Previous single-modality methods cannot be adapted directly for multi-modal data. Table 2 shows the performance of adapted self-supervised federated learning methods which focus on single-modality. From the results, we observe that our method beats these baselines by a large margin, which indicates that previous single-modality methods cannot be easily adapted directly for multi-modal data. Particularly, FedMAE and the FedX perform much worse than other federated learning strategies. We conjecture this is because both of them have ignored modeling the cross-modal alignment. Specifically, FedMAE’s masking strategies might hurt the alignment of image and text modalities, and FedX’s single-modality strategy can’t be adapted to model the image-text cross-modal interaction. Instead, we think FedAvg is still a strong baseline compared to other adapted methods, as we observed in the experiments. That’s because FedAvg doesn’t hurt the ability to learn cross-modal alignment of backbone pre-training models by adding augmented losses. Our method can be transferred to multiple multi-modal pre-training methods. Table 3 shows the downstream task performance of the MGCA and the GLoRIA backbone pre-training method when combined with our strategy. Our method has successfully upgraded MGCA and GLoRIA to adapt to the heterogeneous federated multi-modal pre-training scene, as shown in the significant improvement of the classification and segmentation downstream tasks.
5.4 Analysis experiment
In this section, we mainly conduct analysis experiments using the SLIA backbone method for pre-training unless specifically stated.
Table 4: Ablation results. We report the few-shot classification accuracy and the retrieval accuracy.
Strategy $1\%$ (cls.) Chest imaG. (ret.)
Covid RSNA Rec.@1 Rec.@5 Wst.@1 Wst.@5
[…]
ours 83.21 80.50 29.49 26.99 73.03 67.65
Table 5: Ablation experiment results on the zero-shot generalization ability of our method on the real-world dataset.
Strategy O.O.D In domain In domain Wst.
Rec.@1 Rec.@5 Rec.@1 Rec.@5 Rec.@1 Rec.@5
FedX 12.31 61.04 3.22 16.76 1.58 11.94
FedMAE 12.93 63.72 5.05 23.91 4.34 20.86
FedEMA 15.19 66.94 24.01 67.05 21.85 62.44
FedAvg 15.3 66.62 28.83 72.12 25.34 66.68
FedRGB (Ours) 15.43 67.29 29.49 73.03 26.99 67.65
Figure 8: The trade-off adjusted by the hyper-parameter $\alpha$.
Figure 9: The role of the hyper-parameter $\alpha$.
Figure 10: The role of the radius of the uncertainty set $\rho$.
Robust teacher align module is better than simply averaged one. To investigate the efficiency of the robust align module, we replace it with an align module that is simply averaged by local ones and pre-train a new model. As shown in table 4, the downstream performance of the replaced one is lower than the original one. This means the robust align module can enhance the ability to capture cross-modal alignments and learn transferable representation. Strong regularization is better than weak regularization for guiding the updation. To investigate the effect of the guidance from the robust align module, we replace the L-2 regularization with contrastive loss as a weaker regularization and pre-train a new model. As shown in table 4, the weaker guidance leads to lower classification accuracy and retrieval accuracy in the downstream tasks. Guidance hyper-parameter adjusts the average and disparity of client performances. Figure 8 shows that higher guidance achieves smaller disparity while bringing less harm on average performance. We changed the hyperparameter $\alpha$ from 2 to 10, the disparity of the retrieval accuracy on test sets of different clients decreased significantly while there was only a slight loss on the averaged performance. On the other hand, Figure 9 shows that increasing $\alpha$ wouldn’t result in a penalty in the transferability of the pre-trained backbone, the downstream task performances are maintained at a high level as we enlarge the guidance. Detailed results can be seen in Table 15.
Figure 11: Robustness & convergence on the number of com. turns.
Convergence check: The downstream performance of the pre-trained got promoted as we increased the number of comm. turns. After 20 communications, the performance turned out to be stable which verified the convergent ability of our method. Detailed results can be seen in Table 14. Our method has better classification accuracies on unseen out-of-distribution multi-modal X-Ray datasets. The multi-modal retrieval accuracy in the unseen domain of our method beats the baseline by a large margin as shown in Table 5, which verifies our method’s ability to learn generalizable cross-modal alignment. Meanwhile, our FedRGB has higher in-domain averaged and worst-case acc.. Therefore, improvements in the in-domain performance of FedRGB do not come at the cost of decreased out-of-domain performance. Smaller radius of uncertain set brings higher generalization and transferability on downstream task. Figure 10 shows the downstream performance of models pre-trained with different radius of uncertainty set in the DRO process. As the $\rho$ increases, the performance of fine-tuning-based downstream tasks shows an upward trend. Detailed results can be seen in Table 10.
6 Conclusion
Data limitation is a long-standing problem in the multi-modal learning domain. Despite federated learning can leveraging datasets from multiple sources while guaranteeing privacy issues, its performance would be damaged by data heterogeneity. Inspired by our empirical findings on the impact of heterogeneity on federated multi-modal learning, we propose the FedRGB framework to multigate heterogeneity for federated biomedical vision-language pre-training. The efficiency of our method has been verified by comprehensive experiments.
7 Limitation
The design of the proposed FedRGB approach has taken it into serious consideration to mitigate the potential negative societal impact.. For example, to tackle the data privacy issue, FedRGB has avoided the transmission of any data or latent representations during all server-client communications, while only transferring model parameters and loss, thereby minimizing the risk of privacy leakage in federated learning as much as possible. However, our model may still be subject to privacy-related issues due to the communication of model parameters, which is a common problem faced by other federated learning methods as well. There are also research studies in this line to avoid privacy leakage from model parameters. For fairness, our method aims at optimizing the performance of the worst-case client to promote model equity across different cohorts, whichhas shown great generalizable ability on unseen data. This approach ensures that the pre-trained model not only performs well and fairly on the data from hospitals participating in the federated training but also demonstrates good performance on hospitals that did not participate in the training. Meanwhile we acknowledge that there may still exist limitations. Firstly, If we can resolve the trade-off between privacy protection and data representation communication, it may allow us to design more advanced algorithms to further enhance the performance of federated multimodal pre-training models. Furthermore, in this work we mainly consider the multimodal medical datasets of X-Ray imageeport due to the limitation of the public datasets. However, some rare diseases and more complicated modalities in various medical applications lack paired data, which] may bring other challenges for federated multimodal pretraining.
8 Acknowledgement
Zitao Shuai and Liyue Shen acknowledges support from U-M MIDAS PODS Grant and U-M MICDE Catalyst Grant, and computing resource support from NSF ACCESS Program. This work used NCSA Delta GPU through allocation CIS230133 and ELE230011 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services &\&& Support (ACCESS) program, which is supported by National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296.
Appendix A Problem Formulation
We aim at mitigating the data heterogeneity issues in federally pre-training biomedical vision-language models to better capture the semantic alignment between the two modalities, with heterogeneous private datasets from different clients. In this section, we will formulate our problem set-up, assumptions, and related notations. Federated Vision-Language Pre-training. In this paper, we consider the federated vision-language pre-training (VLP) problem in the biomedical domain. Suppose we are given $N$ clients $\{C_{i}\}_{i=1}^{N}$, each client has a corresponding local yet private dataset $\{D_{i}\}_{i=1}^{N}$ that can’t be easily shared due to privacy issues. Without access to centralized pre-training, federated pre-training aims to conduct vision-language representation learning across local clients, while still capture a robust and generalizable cross-modal alignment, by efficiently aggregating client models $\{M_{i}\}_{i=1}^{N}$ trained on their private datasets. Similar to conventional federated learning scenarios, our federated pre-training consists of multiple communication turns. For communication turn $t$, client models $\{M_{i}\}_{i=1}^{N}$ completed the same number of local iteration steps respectively, and the server receives and aggregates the weights sent from local models to produce the aggregated server model $M^{t}$. Then it will be sent back to clients and re-trained on $\{D_{i}\}_{i=1}^{N}$ separately to get new local models $\{M_{i}^{t+1}\}_{i=1}^{N}$. The federated vision-language pre-training is a more practical and general task compared to related works. Table 1 lists the most similar settings in previous works, and the difference between our task and theirs. Data Formulation. We consider the multimodal datasets including vision and language modalities $X,Y$, consisting of respective data samples $X=\{x_{i}\}_{i=1}^{n},Y=\{y_{i}\}_{i=1}^{m},n,m\in N^{+}$. We assume there exists latent variables U that embed semantics across two modalities such that $X=T_{x}(U),Y=T_{y}(U)$, where $T_{x}(\cdot),T_{y}(\cdot)$ are assumed transformations that is not explicit to know. For example, the disease category of a given X-ray radiology-report pair is a latent variable connecting these two modalities, which determines the pathology region of the radiology image and semantics in the diagnosis report. Thus, we assume each client has paired image-text data that are implicitly correlated through the same latent variables. Therefore, local client datasets could be written as $\mathcal{D}:\{D_{i}:(T_{x}(u),T_{y}(u)),u\sim P_{i}(U)\}_{i=0}^{N}$. Here, we assume the local datasets are heterogeneous, which means the distribution of latent variables $P_{i}(U)$ is different across clients. We formulate two distributions $\mu_{i}\in P(X_{i}),\nu_{i}\in P(Y_{i})$ following (Chen et al., 2020a) as $\mu_{i}=\sum_{j=1}^{n_{i}}a_{i}\delta_{T_{x}(u_{j}))_{i,j}}$ and $\nu_{i}=\sum_{j=1}^{n_{i}}b_{j}\delta_{T_{y}(u_{j})_{i,j}}$, with $\delta_{x}$ as the Dirac function centered on $x$. The weight vectors $a_{i}=\{a_{j}\}_{j=1}^{n_{i}}\in\Delta_{n_{i}}$ and $b_{i}=\{b_{j}\}_{j=1}^{n_{i}}\in\Delta_{n_{i}}$ belong to the $(n-1)$- dim simplex, where $\mu_{i}$ and $\nu_{i}$ are probability distributions. This formulation can also be applied to feature vectors or representations. For a given set of distributions $\{P_{i}\}_{i=1}^{N}$, we can form a new distribution $P^{c}=\sum_{i=1}^{N}w_{i}P_{i},\textbf{w}\in\Delta$ where $\Delta$ is a $(N-1)$-dim simplex by weighted averaging the sample vectors. Model Architectures. Following the VLP works (Radford et al., 2021; Yang et al., 2022), we frame the VLP using a dual-encoder structure, where the model for each modality consists of a feature encoder module $f(\cdot)$ that outputs feature vectors given the raw input, and a feature alignment module $g(\cdot)$ that projects feature vectors to aligned representations in the shared embedding space. Denote the feature vector as $s$, and the aligned representation as $z$. We model the outputs of the encoder and alignment modules as conditional distributions. Specifically, the encoder takes the multi-modal raw input to generate semantic embeddings of the corresponding modality, which could be formulated as $P(s|U)$. Similarly, that of the alignment module can be formulated as $P(z|S,U)$.
Appendix B Details of Learning Unbiased Cross-modal Alignment via Distribution-based Optimization
In order to guide the local client training, we need to provide a teacher alignment module as mentioned in the last section, which can capture globally unbiased and generalizable cross-domain alignment. How can we obtain such a teacher guidance model? We propose to learn the unbiased cross-domain alignment by considering the robustness across various distributions of local clients. A teacher alignment module should avoid distorting the aggregated feature encoder while learning the cross-modal alignments. Additionally, we introduce a much harder task to force the module to be more robust to biased alignment, by considering the influence of the biased alignment module on the feature encoder. Furthermore, to learn a teacher model that performs well on varies data distributions, we minimize distortion degree under the most biased local alignment with the “worst-case” local data distribution. Inspired by the distributionally robust optimization (DRO) (Cotter et al., 2019), we model the de-centralized pre-training framework as a two-stage DRO learning process. In the first stage, the local models are trained with $\mathcal{L}_{local}$ (Eq. 2), while the whole model is jointly optimized with the guidance of the teacher alignment module. In the second stage, we learn a robust teacher alignment module that performs well conditioned on the worst-case local models where the feature encoder is distorted the most severely. We aim to optimize the alignment module on the worst-case distribution of an uncertainty set constructed by the local aligner and local data distribution, which represents possible cases that the model would encounter. We construct the uncertainty set based on the distribution of client datasets and local aligners learned in the first stage. From the probabilistic view, the local aligners could be written as $P(z|U,f)$, which takes the output of the given encoder $f$ to generate the aligned representations, while $f$ takes the raw input from the local data distribution $P(U)$ to generate the semantic embedding that has not well learned the cross-modal alignment. Therefore, the distribution to be considered should be $P(z,U|f)$. Hence the uncertainty set can be written as:
$\footnotesize Q^{c}:\{D(Q||P^{c})\leq\rho\},P^{c}=\sum_{i=1}^{N}w_{i}P_{i}(z,U|f),w\in\Delta_{N-1}$ (7)
where $N$ is the number of clients, $\Delta$ is a $(N-1)$-dim simplex, $D$ is a discrepancy measure, and $\rho$ is the uncertainty radius that measures the uncertainty degree of the unseen distributions and influences model’s generalization ability. Finally, with the teacher align module denoted as $g^{*}$, indicating the aggregated encoder after the first stage training, the learning object in the second stage can be formulated as:
$\footnotesize min_{g^{*}}\ max_{P\in Q^{c}}\ \mathcal{L}_{dro}(g^{*},P,f^{*})$ (8)
Following (Zhang et al., 2023b), the optimization goal can be constructed by distributions $P_{i}$ of the support set:
$\footnotesize min_{g^{*}}\ max_{\textbf{w}\in\Delta_{N-1}}\sum_{i=1}^{N}w_{i}\ \mathcal{L}_{dro}(g^{*},P_{i},f),s.t.D(N\textbf{w}||\textbf{1})\leq\rho$ (9)
where $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$ are the losses calculated on the local clients respectively. While we learn $g^{*}$ by gradient descent, the mirror gradient descent process of optimizing w could be solved efficiently (Zhang et al., 2023b; Duchi et al., 2008):
$\footnotesize{w_{i}}=\text{Proj}_{\mathcal{Q}{c}}\left({w_{i}e^{\gamma v_{i}}}/\sum_{i=1}^{N}{w_{i}e^{\gamma v_{i}}},\rho\right)$ (10)
where $\text{Proj}_{\mathcal{Q}{c}}$ refers to the projection to $\{w:D(N\textbf{w}||\textbf{1})\leq\rho\}$, $v_{i}$ represents the loss $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$. We alternatively optimize the w and the teacher align module. For each communication turn of the federated learning, we first obtain an averaged encoder, and a set of local align modules in the first stage. Then we solve the weights and optimize the parameters of the teacher align module. Hence the w could be written in a more trackable form as:
$\footnotesize{w_{i}^{(r+1)}}=\text{Proj}{\mathcal{Q}{c}}\left({w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}}/\sum_{i=1}^{N}{w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}},\rho\right)$ (11)
where $r$ refers to the communication turn.. Finally, we formulate $\mathcal{L}_{dro}$ to measure the performance of the teacher align module of alleviating distortion on the encoder. Based on equations (3) and (4), we update the teacher alignment module in local clients with weights w. Specifically, given a client $c$, we use $f^{*}$ learned in the first stage of the current communication turn to generate features $(\tilde{v^{c}},\tilde{l^{c}})$, and a learnable copy of $f^{*}$ to generate distorted features $(v^{c},l^{c})$. The distorted features are input to the frozen $g$ to calculate the task loss $L_{task}$, and are input to the $g^{*}$ to get calibrated representations $(z_{I}^{c*},z_{T}^{c*})$. Meanwhile, $(\tilde{v^{c}},\tilde{l^{c}})$ are input to $g^{*}$ to get representation $\tilde{(z_{I}^{c*}},\tilde{z_{T}^{c*}})$ for regularization. The regularization term could be written as:
$\footnotesize\mathcal{L}_{dro,\tau}=\beta(\sum_{i=1}^{N_{c}}(\tilde{z^{c*}}_{I,i}-z^{c*}_{I,i})^{2}+\sum_{i=1}^{N_{c}}(\tilde{z^{c*}}_{T,i}-z^{c*}_{T,i})^{2})+\mathcal{L}_{CL,\tau}$ (12)
where $\mathcal{L}_{CL}$ is the contrastive loss (Radford et al., 2021) between $(z^{fz,c*}_{img,i},z^{fz,c*}_{txt,i})_{i=0}^{N_{c}}$, and $N_{c}$ refers to the number of samples in the client $c$.
Appendix C Experiment Details
C.1 Pre-training setup
Following (Wang et al., 2022) we utilize the MIMIC-CXR (Bigolin Lanfredi et al., 2022) dataset for multi-modal pre-training. This dataset is widely used in the biomedical multi-modal learning domain, with $227,835$ image-text pairs from $65,379$ patients. Some related works also have imported additional features to image-text pairs to augment the data. However, we only use the image-text pairs for pre-training to make the results more generalizable. During the pre-training, local clients only have access to their highly heterogeneous datasets. To construct the heterogeneous client datasets, following (Yan et al., 2023) we employ the Latent Dirichlet Allocation (LDA) (Blei et al., 2003) to divide the MIMIC-CXR dataset into $5$ partitions based on a selected sensitive attribute. For implementation, we import the corresponding attribute information of given image-text pairs from the Chest-Imagenone dataset (Wu et al., 2021) and divide local datasets based on disease category. The disease category is a multi-label binary attribute and is transformed into a multi-class label. That’s because the words in the clinical report are highly related to the disease category as illustrated in 12. We set the heterogeneity degree in the LDA algorithm to be 1 and the subgroups of the MIMIC-CXR that are used as local datasets are summarized in Table 7. Specifically, we select 5 commonly considered diseases (Bannur et al., 2023): ’Edema’,’Pleural Effusion’, ’Consolidation’, ’Pneumothorax’, and ’Pneumonia’. We set the non-NaN value to 1 and then set NaN value to 0 to construct a 5-way binary multi-label. Then we get $2^{5}$-category multi-class label, and we run LDA on it. The distribution of the disease of the client dataset is summarized in Table 7.
Figure 12: Illustration of the divided heterogeneous subgroups for each client.
Table 6: Distortion of local training of FedAvg method. We show the performance gap between the server and local retrained models. We employ the decrease in the similarity of the image embedding and text embedding to be an indicator of distortion.
FedAvg Test on C1 Test on C2 Test on C3 Test on C4 Test on C5
Retrain On C1 0.092 0.078 0.08 0.073 0.082
Retrain On C2 0.054 0.049 0.046 0.047 0.048
Retrain On C3 0.069 0.072 0.07 0.066 0.067
Retrain On C4 0.092 0.09 0.084 0.093 0.088
Retrain On C5 0.073 0.08 0.07 0.068 0.087
Avg 0.076 0.0738 0.07 0.0694 0.0744
Table 7: The distribution of the sensitive attribute age-decile. From this table, we can verify that the local datasets are highly heterogeneous.
Edema P. E. Consolid. Pneumothorax Pneumonia C1 C2 C3 C4 C5
8251 52931 23153 16054 2983
Yes 367 130 4362 1993 2671
Yes 67 11425 5825 778 655
Yes Yes 945 1693 3472 2640 1607
Yes 1641 261 599 791 1057
Yes Yes 420 61 198 245 14
Yes Yes 253 50 18 805 913
Yes Yes Yes 434 95 27 303 153
Yes 476 2504 128 8791 1097
Yes Yes 827 284 559 603 156
Yes Yes 4366 1866 3507 2815 1140
Yes Yes Yes 5965 217 510 109 921
Yes Yes 123 80 444 111 84
Yes Yes Yes 48 9 101 4 269
Yes Yes Yes 807 121 3 922 670
Yes Yes Yes Yes 723 818 42 246 69
Yes 7447 3656 1505 246 1395
Yes Yes 278 927 719 1547 453
Yes Yes 98 970 $1$ 639 1145
Yes Yes Yes 2777 3019 2276 105 466
Yes Yes 10 323 120 227 760
Yes Yes Yes 31 57 153 28 9
Yes Yes Yes 30 16 238 148 91
Yes Yes Yes Yes 92 a74 9 101 36
Yes Yes 208 286 48 15 33
Yes Yes Yes 70 214 248 1 90
Yes Yes Yes 164 76 39 126 872
Yes Yes Yes Yes 133 251 417 155 1224
Yes Yes Yes 71 6 15 8 8
Yes Yes Yes Yes 2 56 42 3 20
Yes Yes Yes Yes 156 18 11 44 279
Yes Yes Yes Yes Yes 38 156 201 38 183
We divide the MIMIC-CXR into 5 heterogeneous subgroups to construct 5 client datasets. Each divided dataset consists of train splits and validation splits based on the notation of the MIMIC-CXR. We use $404\times A$ 40 for pre-training and set the batch size to 388. We set the learning rate to $2\times 10^{-5}$, the number of communications to 25. We set the updating step size $\gamma=1$, and uncertainty radius $\rho=1$, hyper-parameter $\alpha=1,\beta=2$ in the main experiments. For each communication, we randomly sample 50 batches of data from the client datasets. To obtain the guidance, we run an extra communication for our method.
C.2 Downstream tasks
We conduct Few-shot classification, Medical image segmentation, and Image Retrieval downstream tasks to evaluate the generalization ability of the pre-trained model. Few-shot classification. To verify the efficiency of the pre-trained model on the general medical image task, we test their performance on multiple image classification benchmarks: (1) RSNA Pneumonia Detection (RSNA) (Shih et al., 2019), the task is to predict if the image reflects a pneumonia-positive phenomenon. (2) Covidx (Wang et al., 2020), which has three categories that COVID-19, non-COVID pneumonia and normal. We fine-tune our pre-trained model with an additional linear layer on 1%, 10% percentage of the training dataset of downstream tasks, and evaluate the image classification accuracy. For evaluation, we report the classification accuracy of the models on them. Medical image segmentation. To investigate the transfer ability of our model on fine-grained tasks, we conduct experiments on medical image segmentation on the RSNA (Wang et al., 2020) benchmark. We follow (Wang et al., 2022) to convert object detection ground truths of RSNA into masks for semantic segmentation. Similar to (Huang et al., 2021), we construct a U-Net framework using a decoder and our pre-trained image encoder. During the fine-tuning, we freeze the encoder and fine-tune the decoder using 1%, 10% of the training data of the downstream dataset, and then use the Dice score to evaluate the performance of our model. Image retrieval. To verify if the pre-trained models have learned the semantics and alignment relationships between the image and text contained in the pre-training dataset, we conduct the image retrieval downstream task. Specifically, we utilize the validation splits of the local clients to test the image retrieval performance of the aggregated models. For a text from a given batch of input image-text pairs, we calculate the similarities between this text and images in this batch. Then we sort these similarities and select the top-1 and top-5 images from them. If the corresponding image of the text is contained in the selected set, then the image is retrieved correctly. Here we utilize the top-1 recall accuracy and top-5 recall accuracy to evaluate the pre-trained models. Out-of-Distribution testing benchmark. We adapt the EHRXQA (Bae et al., 2023) VQA dataset for evaluation. This dataset is built on the meta-data notation of MIMIC-CXR (Bigolin Lanfredi et al., 2022) and ChestImaGenome (Wu et al., 2021). We merge the question and answer sentences as the text input and use the corresponding image to construct the image-text pairs for the multi-modal retrieval task. Other settings are the same as the ordinary image retrieval task.
C.3 Backbones and Baselines
We focus on upgrading biomedical vision-language pre-training (VLP) strategies to heterogeneous federated learning scenes. Since federated heterogeneous VLP is a new topic, we adopt existing biomedical VLP baselines to federated strategies (Lu et al., 2023a). Specifically, we consider adopting simple language-image contrastive alignment (ConVIRT) (Zhang et al., 2022b; Radford et al., 2021), global-local language-image contrastive alignment(GLoRIA) (Huang et al., 2021), and Multi-Granularity Cross-modal Alignment (MGCA) (Wang et al., 2022). For federated learning strategies, we consider simple averaging (FedAvg) (McMahan et al., 2017; Lu et al., 2023a), centralized training (Local), and decentralized training (Global). For fair comparisons, we combine each federated learning strategy with different biomedical VLP methods. For baselines pre-trained in Local strategy, we report the averaged performance of the local models. For fair comparisons, we implement the method with the same backbones as our method. We utilize a ViT-Base model as our vision encoder and utilize the Bert-Base as our text encoder. We pre-process the input image following (Wang et al., 2022) to resize the radiology graph to the size of $224\times 224$ to match the vision encoder.
Table 8: Descriptive Statistics of the subgroups.
Disease C1 C2 C3 C4 C5
Edema 13150 8061 13336 8121 8341
P.E. 17048 20865 17706 9974 10424
Consolid. 4879 2201 2221 4024 4615
Pneumoth. 14177 6962 6315 13991 7115
Pneumonia 11605 10105 7152 3431 7064
Appendix D Additional Analysis Experiment
We report the communication costs and training costs of our method and other FL methods on both the server and client sides in Table 9. As shown in the table, the additional computational overhead introduced by our method is within an acceptable range.
Table 9: Summarization of computation costs.
Method Train Params Train time Server Cost Info. Trans.
De-centralized 172.6 M 14.62s/it 863M FLOP 172.6 M
FedAvg 172.6 M 14.62s/it 863M FLOP 172.6 M
FedMAE 226.1M 14.95s/it 1130.5M FLOP 226.1M
FedX 172.6 M 22.66s/it 863M FLOP 172.6 M
FedEMA 172.6 M 14.59s/it 863M FLOP 172.6 M
Ours [1st stage] 190.4M 15.4s/it 951.2M FLOP 190.4M
Ours [2nd stage] 190.4M 23.68s/it 89M FLOP 17.8M
Centralized 172.6 M 14.62s/it 863M FLOP 172.6 M
Table 10: Detailed results on the analysis experiments of the role of hyper-parameter $\alpha$.
$\rho$ RSNA (cls.) Covid (cls.) RSNA (seg.)
$1\%$ $10\%$ $1\%$ $10\%$ $1\%$ $10\%$
$\rho=1e-7$ 83.51 84.04 81.00 91.00 70.44 73.05
$\rho=1e-4$ 83.51 83.54 82.25 90.00 70.39 73.11
$\rho=1$ 83.16 83.41 80.00 89.75 66.98 70.18
Our method has succesfully avoided the distortion caused by heterogeneous datasets during the local training. Empirically, we demonstrate the effectiveness of this guidance-based local training strategy. Specifically, we retrain the model on each local dataset with the same backbone methods and setting and conduct the study on the change of similarity between embeddings of paired image and text. We show the performance gap between the server and local retrained models. The reduced similarity of the image embedding and text embedding from paired data can be regarded as an indicator of ”distortion” of cross-modal alignment. As reported in Table 6 and Table 11, we observe that the average decrease on the similarity of image and text embedding of our method is much smaller compared to FedAvg baseline. Role of the number of clients that participated in the federated pre-training. We have conducted experiments on how the number of clients participating in the federated pre-training affects the performance of the learned model. As shown in the Table 16, we have observed that bringing more diverse clients for the federated pre-training is likely to improve the performance of the pre-trained model.
Appendix E Detailed Experiment Results
Here we show the detailed experiment results.
Table 11: Distortion of local training of ours method. We show the performance gap between the server and local retrained models. We employ the decrease in the similarity of the image embedding and text embedding to be an indicator of distortion.
Ours Test on C1 Test on C2 Test on C3 Test on C4 Test on C5
Retrain On C1 0.025 0.023 0.017 0.034 0.01
Retrain On C2 0.046 0.05 0.043 0.057 0.063
Retrain On C3 0.05 0.052 0.044 0.055 0.036
Retrain On C4 0.062 0.043 0.038 0.041 0.033
Retrain On C5 0.071 0.058 0.048 0.066 0.053
Avg 0.0508 0.0452 0.038 0.0506 0.039
[…]
Strategy Recall@1 (ACC) Recall@5 (ACC)
C1 C2 C3 C4 C5 Avg. C1 C2 C3 C4 C5 Avg.
Centralized 43.65 38.55 40.08 43.10 41.86 44.45 86.63 79.98 82.65 84.99 86.83 84.22
FedAvg 30.43 25.34 26.90 28.75 32.73 28.83 76.39 66.68 69.75 73.84 73.94 72.12
[…]
Local.avg. 17.73 13.62 14.32 15.83 16.04 15.51 55.43 45.97 48.30 51.38 51.12 51.12
[…]
Strategy com. turn Recall@1 (ACC) Recall@5 (ACC)
C1 C2 C3 C4 C5 C1 C2 C3 C4 C5
FedAvg 25 30.43 25.34 26.90 28.75 32.73 76.39 66.68 69.75 73.84 73.94
[…]
1-5 Avg. 25 29.19$\downarrow$ 23.45$\downarrow$ 24.68$\downarrow$ 26.07$\downarrow$ 29.14$\downarrow$ 73.20$\downarrow$ 64.40$\downarrow$ 66.89$\downarrow$ 70.20$\downarrow$ 69.79$\downarrow$
[…]
1-5 Avg. 26 29.32$\downarrow$ 23.51$\downarrow$ 24.70$\downarrow$ 26.26$\downarrow$ 29.38$\downarrow$ 73.33$\downarrow$ 64.44$\downarrow$ 66.34$\downarrow$ 70.10$\downarrow$ 69.41$\downarrow$
Table 14: Robustness check on communication turns.
Com. RSNA (cls.) Covid (cls.) RSNA (seg.) Chest imaG. (ret.)
$1\%$ $10\%$ $1\%$ $10\%$ $1\%$ $10\%$ Rec.@1 Rec.@5 Wst.@1 Wst.@5
commu.=0 79.89 82.11 71.50 85.50 62.03 67.46 3.27 16.44 2.53 13.54
commu.=15 83.16 83.41 80.00 89.75 66.98 70.18 28.56 71.83 2595 67.21
commu.=20 83.26 83.39 80.00 90.25 69.93 70.76 29.69 72.09 26.75 67.52
commu.=25 83.34 83.39 80.75 91.25 71.42 72.89 29.80 72.59 27.48 68.64
Table 15: Detailed results on the analysis experiments of the role of hyper-parameter $\alpha$.
Hyper. Alpha RSNA (cls.) Covid (cls.) RSNA (seg.) Chest imaG. (ret.)
$1\%$ $10\%$ $1\%$ $10\%$ $1\%$ $10\%$ Recall Acc. Worst Acc. Disparity
$\alpha=$2 83.49 83.83 79.99 91.00 70.67 72.27 3.25 28.91 25.95
$\alpha=$5 83.51 84.01 79.50 90.00 70.31 72.44 2.99 28.58 26.12
$\alpha=$10 83.16 83.41 80.00 89.75 66.98 70.18 2.37 28.56 26.17
Table 16: Detailed results on the analysis experiments of the effect of the number of participated clients.
Num. Client RSNA (cls.) Covid (cls.) RSNA (seg.) Chest imaG. (ret.)
$1\%$ $10\%$ $1\%$ $10\%$ $1\%$ $10\%$ Rec@1 Acc. Wst@1 Acc. Rec@5 Acc. Wst@5 Acc.
n=2 82.11 83.06 77.25 88.5 62.73 71.17 22.27 62.2 18.7 56.23
n=5 83.21 83.54 80.5 90 71.51 74.08 29.49 73.03 26.99 67.65
[…]
position model com. Recall@1 (ACC) Recall@5 (ACC)
C0 C1 C2 C3 C4 Avg. C0 C1 C2 C3 C4 Avg.
- server 25 30.43 25.34 26.90 28.75 32.73 28.83 76.39 66.68 69.75 73.84 73.94 72.12
- server 50 32.29 26.03 26.99 27.06 30.23 28.52$\downarrow$ 77.60 67.86 69.43 72.06 71.74 71.74$\downarrow$
[…]
strategy model Recall@1 (ACC) Recall@5 (ACC)
C1 C2 C3 C4 C5 Avg. C1 C2 C3 C4 C5 Avg.
Global server 43.65 38.55 40.08 43.10 41.86 41.45 86.63 79.98 82.65 84.99 86.83 84.22
FedAvg server 30.43 25.34 26.90 28.75 32.73 28.83 76.39 66.68 69.75 73.84 73.94 72.12
[…]
Table 19: Detailed ablation results.
Strategy Recall@1 (ACC) Recall@5 (ACC)
C1 C2 C3 C4 C5 Avg. C1 C2 C3 C4 C5 Avg.
[…]
ours 32.53 26.99 27.70 29.09 31.12 29.62 77.08 67.65 70.32 73.96 76.12 73.14

### v2

@@
-Mitigating Heterogeneity in Federated Multimodal Learning with Biomedical Vision-Language Pre-training
+Align as Ideal: Cross-Modal Alignment Binding for Federated Medical Vision-Language Pre-training
 Abstract
-abstract Vision-language pre-training (VLP) has arised as an efficient scheme for multimodal representation learning, but it requires large-scale multimodal data for pre-training, making it an obstacle especially for biomedical applications.
-To overcome the data limitation, federated learning (FL) can be a promising strategy to scale up the dataset for biomedical VLP while protecting data privacy.
+Vision-language pre-training (VLP) has arised as an efficient scheme for multimodal representation learning, but it requires large-scale multimodal data for pre-training, making it an obstacle especially for medical applications.
+To overcome the data limitation, federated learning (FL) can be a promising strategy to scale up the dataset for medical VLP while protecting data privacy.
 However, client data are often heterogeneous in real-world scenarios, and we observe that local training on heterogeneous client data would distort the multimodal representation learning and lead to biased cross-modal alignment.
-To address this challenge, we propose Federated distributional Robust Guidance-Based (FedRGB) learning framework for federated VLP with robustness to data heterogeneity.
-Specifically, we utilize a guidance-based local training scheme to reduce feature distortions, and employ a distribution-based min-max optimization to learn unbiased cross-modal alignment.
-The experiments on real-world datasets show our method successfully promotes efficient federated multimodal learning for biomedical VLP with data heterogeneity.
-Vision-Language Pre-training, Federated Learning, Data Heterogeneity
-abstract Vision-language pre-training (VLP) has arised as an efficient scheme for multimodal representation learning, but it requires large-scale multimodal data for pre-training, making it an obstacle especially for biomedical applications.
-To overcome the data limitation, federated learning (FL) can be a promising strategy to scale up the dataset for biomedical VLP while protecting data privacy.
-However, client data are often heterogeneous in real-world scenarios, and we observe that local training on heterogeneous client data would distort the multimodal representation learning and lead to biased cross-modal alignment.
-To address this challenge, we propose Federated distributional Robust Guidance-Based (FedRGB) learning framework for federated VLP with robustness to data heterogeneity.
-Specifically, we utilize a guidance-based local training scheme to reduce feature distortions, and employ a distribution-based min-max optimization to learn unbiased cross-modal alignment.
-The experiments on real-world datasets show our method successfully promotes efficient federated multimodal learning for biomedical VLP with data heterogeneity.
-Figure 1: Federated multi-modal pre-training aims to de-centralizedly utilizing private dataset to tackle the data hungry problem.
-Fitting heterogeneous local datasets that correspond to different distributions of latent semantics (e.g.
-disease) would result in biased models.
-Our FedRGB framework has successfully upgraded multi-modal pretraining methods by introducing robust alignment module as guidance during local training time.
+To address this challenge, we propose a Federated Align as IDeal (FedAID) framework for federated VLP with robustness to data heterogeneity, to bind local clients with an ideal cross-modal alignment.
+Specifically, to reduce distortions on global-aggregated features while learning diverse semantics from client datasets during local training, we propose to bind the cross-model aligned representation space learned by local models with an unbiased one via guidance-based regularization.
+Moreover, we employ a distribution-based min-max optimization to learn the unbiased cross-modal alignment at each communication turn of federated pre-training.
+The experiments on real-world datasets demonstrate our method successfully promotes efficient federated multimodal learning for medical VLP with data heterogeneity.
 1 Introduction
 Vision-language pre-training (VLP) learns efficient multimodal representations by extracting latent semantics from a large amount of image-text pairs, where the dataset scale largely impacts the performance of the learned model (Oquab et al., 2023).
-However, enlarging pre-training dataset is a non-trivial challenge especially for biomedical applications which is restricted by privacy issues of clinical and patient data (Ladbury et al., 2023).
-To utilize data across different institutes, recent work (Lu et al., 2023a) has attempted to introduce federated learning into the biomedical VLP domain.
-However, in real-world scenarios, datasets collected from different institutes are always heterogeneous, which is a long-standing practical challenge in federated learning (Ghosh et al., 2019; Zhang et al., 2023a; Huang et al., 2022).
-For example, one hospital receives a high proportion of patients with pneumonia, while another hospital may undertake more patients with tuberculosis, due to variety of patients cohort and geographics.
-While the data heterogeneity would essentially damage the performance of the federally trained model, current FL methods for biomedical VLP (Lu et al., 2023a) haven’t taken this vital factor into consideration to provide an effective solution.
-To address this issue, simply averaging client models trained on heterogeneous local datasets results in biases as shown in Figure 1.
-Firstly, the client model would be heterogeneous, and re-training on the sent-back server model might distort the aggregated structure of the representation space.
-Secondly, averaging model parameters cannot produce a robust alignment between the two modalities, which not only affects the generalization ability, but also prevents the model from learning more diverse correlations between vision and language.
-Furthermore, as we will demonstrate in our empirical findings (Sec 5.2), heterogeneity harms the performance of pre-trained model by affecting the deep layers which serve as an alignment module in the VLP model.
-In this paper, we propose the FedRGB framework to mitigate heterogeneity in federated biomedical VLP to enable its deployment in real-world scenarios.
-To alleviate the distortion in local client training, we introduce a teacher alignment module to provide unbiased cross-modal alignment by generating anchor representation space for regularization.
-We employ a distributionally robust optimization (DRO)-based algorithm to learn a robust teacher alignment module, which is optimized to better alleviate the worst-case distortion on the feature encoder caused by local cross-modal alignment.
+However, enlarging multimodal pre-training dataset is a non-trivial challenge especially for medical applications which is restricted by privacy issues of clinical and patient data Ladbury et al.
+(2023).
+To collect and utilize data across different institutes, recent work Lu et al.
+(2023a) has attempted to introduce federated learning into the medical VLP domain.
+However, in real-world scenarios, datasets collected from different institutes are always heterogeneous, which is a long-standing practical challenge in federated learning Ghosh et al.
+(2019); Zhang et al.
+(2023a); Huang et al.
+(2022).
+For example, one hospital receives a high proportion of patients with pneumonia, while another hospital may undertake more patients with Edema, due to the variety of patient cohorts and geographics.
+While the data heterogeneity would essentially damage the performance of the federally trained model, current medical multi-modal pre-training methods haven’t taken this vital factor into consideration, which encounters the deployment of medical VLP into decentralized pre-training scenarios.
+For a thorough study especially under the case of multi-modal learning, while multi-modal pre-training aims to learn a well-aligned latent space shared by each modality Zhang et al.
+(2022b); Li et al.
+(2021a), this cross-modal alignment would be affected by the data heterogeneity issue during local training Zhang et al.
+(2021); Ye et al.
+(2023), thus failing to provide an effective solution.
+The current solution of simply averaging client models trained on heterogeneous local datasets is sub-optimal as shown in Fig.
+1.
+First of all, when the client datasets are heterogeneous, re-training on the sent-back server model might distort the aggregated structure of the representation space.
+Secondly, as observed in empirical study Sec 5.2, averaging model parameters cannot produce an ideal alignment between two modalities which captures robust and factual diverse correlations between vision and language.
+And biased cross-modal alignment which overfits multi-modal correlations of local datasets, will not only affect generalization ability, but prevent the model from learning transferable and diverse semantics from local datasets.
+Furthermore, as we demonstrate in our empirical findings (Sec 5.2), heterogeneity harms the performance of pre-trained models by affecting the deep layers, which serve as alignment modules in the VLP model.
+Figure 1: Federated multi-modal pre-training aims to decentralizedly utilize private datasets to tackle the data-hungry problem in pre-training transferable models.
+During local training, directly fitting heterogeneous local datasets that w.r.t.
+different distributions of latent semantics (e.g.
+disease) would result in non-ideal cross-modal alignment.
+Our method upgrades multi-modal pretraining methods to federated setting by binding local cross-modal alignment to an ideal one during local pre-training.
+In this paper, we propose Federated Align as IDeal (FedAID) framework to mitigate data heterogeneity effect on encountering learning ideal cross-modal alignment in federated vision-language pre-training and enable its deployment in real-world scenarios.
+To alleviate the distortion on global-aggregated features during local client training while still encouraging models to extract diverse semantics from local datasets, we bind the cross-modal aligned representation space learned by local models to an unbiased representation space.
+Specifically, we introduce a teacher alignment module to provide robust cross-modal alignment by generating anchor representation space for regularization.
+We employ a distributionally robust optimization (DRO)-based algorithm to learn a robust teacher alignment module, which is optimized to better alleviate the worst-case distortion on the feature encoder caused by locally heterogeneous cross-modal alignment.
 Our contributions primarily focus on:
-  • We for the first time tackle the problem of biomedical VLP under the federated setting by utilizing heterogeneous datasets from different institutes.
-We conduct empirical studies to analyze the influence of data heterogeneity on federated multimodal learning.
-  • We propose FedRGB framework to mitigate the heterogeneity in federated multimodal learning.
-Specifically, our method introduces the guidance-based local training scheme for learning unbiased cross-modal alignment to alleviate representaion distortions.
-  • Experiment results show the effectiveness of our method in learning multimodal representations under the federated setting by appling to various downstream tasks including image-text retrieval, classification, and segmentation.
+  • We for the first time tackle the problem of medical VLP under the federated setting by utilizing heterogeneous multi-modal datasets from different institutes.
+We conduct empirical studies to analyze the influence of data heterogeneity on federated multi-modal learning.
+  • We propose FedAID to mitigate the data heterogeneity on encountering learning ideal cross-modal alignment in federated VLP.
+Specifically, our method binds representations aligned by client models with unbiased representation space by introducing the guidance-based local training for learning unbiased cross-modal alignment to alleviate feature distortions.
+  • Experiment results show the effectiveness of our method in learning multi-modal representations under the federated setting by appling to various downstream tasks including image-text retrieval, classification, and segmentation.
 2 Related Work
-2.1 Vision-Language Pre-training
-Vision-Language Pre-training (VLP): Leveraging data of different modalities to enable multimodal reasoning and representation learning for various cross-modal tasks has become an active research field in machine learning (Radford et al., 2021).
-Pre-training vision-language models on large datasets and then transferring learned knowledge to downstream tasks has become a popular approach to extract semantics from unlabeled image-text paired datasets (Li et al., 2022a; Bao et al., 2022).
-To find the shared latent space, current works often focus on learning generalizable and robust alignment relationships between the two modalities (Radford et al., 2021; Chen et al., 2020a), which measures the ability to connect and retrieve the representation of the paired data.
-Biomedical Vision-Language Pre-training: Most of the existing works in biomedical vision-language pre-training are based on the twin modality-specific encoders for medical images and reports respectively, with the image-report contrastive loss (Zhang et al., 2022b) to align visual and language representations on a joint latent space.
-A significant improvement has been achieved using the local contrastive loss which aims to align fine-grained representations (Huang et al., 2021).
-Incorporating medical domain knowledge, Wang et al.
-(2022) proposes a well-designed method that aligns the image and report embeddings in three different granularities.
-Recent works have also tried to address other practical issues of misalignments (Bannur et al., 2023).
-However, most of these methods haven’t considered the adaptation to de-centralized scenarios during pre-training and thus might fail to handle the potential damage from data heterogeneity.
-2.2 Federated Learning
-Heterogeneous Federated Learning: Federated learning aims to de-centralize model training by aggregating models trained on local clients to obtain more generalizable models with scaled-up training datasets while protecting data privacy (McMahan et al., 2017), which achieved success in many applications especially those with data privacy issues like biomedical data.
-However, client datasets are heterogeneous in most cases, which damages the performance of the aggregated model (Ghosh et al., 2019).
-Existing methods have tried to solve the data heterogeneity problem in many ways (Li & Wang, 2019; Collins et al., 2021).
-However, most methods either require additional communications on representations (Zhang et al., 2023a) which are not privacy-preserved and inefficient, or need large-scale public datasets for server aggregation (Huang et al., 2022; Yu et al., 2023b) that does not fit the case for biomedical data.
-Besides, remaining methods are often designed for supervised tasks that focus on label distribution (Diao et al., 2020) or feature distribution (Li et al., 2022b), which cannot be directly adapted to the vision-language pre-training task.
-While the self-supervised pre-trained models can be biased due to the heterogeneous datasets in the biomedical domain (Krishnan et al., 2022) and over-reliance on spurious features (Saab et al., 2022), distributionally robust optimization (Deng et al., 2020) can optimize the worst-case performance thus alleviating these issues (Liu et al., 2022).
-Recent works have considered its application in unsupervised and self-supervised learning senarios (Han et al., 2023; Zhang et al., 2022a; Capitani et al., 2024).
-Federated Multi-modal Learning: Federated learning strategies have been extended to multimodal datasets in recent works (Zhou & Wang, 2022), with the goal to reduce training cost and promote personalized deployment of multimodal foundation models (Su et al., 2022; Chen et al., 2023b; Lu et al., 2023b).
-For example, Yu et al.
-(2023a) considers a client-server collaborative pre-training setting, benefiting the pre-training process of the server model by leveraging representations from local modalities.
-Nevertheless, these works require a pre-training stage in the server while ignoring that only on client datasets with limited samples.
-The closest work to ours is Lu et al.
-(2023a), which adopts federated learning to scale up the pre-training dataset.
-However, this work lacks generalization ability and hasn’t considered the data heterogeneity issue that is very common and significant in practice.
+Medical Vision-Language Pre-training.
+Pre-training vision-language models on large-scale datasets and then transferring learned knowledge to downstream tasks has become a popular approach to leverage diverse semantics contained in multi-modal unlabeled data  Li et al.
+(2022a); Bao et al.
+(2022); Radford et al.
+(2021).
+To obtain more generalizable representations, current works aim to learn the latent space shared by each modality using contrastive learning via cross-modal alignment  Radford et al.
+(2021); Chen et al.
+(2020) or multi-modal fusion Li et al.
+(2021a); Chen et al.
+(2022), which connect the representations from the paired data.
+Existing works in medical vision-language model pre-training employ twin modality-specific encoders for mapping medical images and reports to the latent space respectively, and use the image-text contrastive loss Zhang et al.
+(2022b) to align visual and language representations.
+Improvement has been achieved by using the word-patch contrastive loss which aims to align fine-grained cross-modal representations Huang et al.
+(2021a).
+Furthermore, incorporating medical domain knowledge,  Wang et al.
+(2022) proposes to align the image and report embeddings in three different granularities.
+Recent works have also tried to address other practical issues of misalignments Bannur et al.
+(2023).
+However, almost all of the current methods still rely on large-scale pre-training datasets, which impede the adoption to modalities with limited training samples and deployment in real-world scenarios.
+Federated multi-modal learning.
+Federated multi-modal learning Zhou & Wang (2022) aims to facilitate diverse semantics in local datasets in a decentralized and privacy-preserved wayRelated works often focus on leveraging client-side information to train personalized information while utilizing server-side pre-trained models.
+Su et al.
+(2022); Chen et al.
+(2023); Lu et al.
+(2023b), and use local datasets to enhance the centralized training Yu et al.
+(2023).
+However, these works haven’t answered a key question: how can we decentralize the multi-modal pre-training through a federated way in real-world scenario to learn generalizable cross-modal representations?
+Recent advances Lu et al.
+(2023a) has verified that federated learning can be utilized to scale up the pre-training dataset.
+However, this work hasn’t considered the harm of data heterogeneity issue Ghosh et al.
+(2019), which is very common and significant in practice such as in medical applications.
+Heterogeneity in federated learning.
+Existing methods have tried to solve this problem in uni-modal supervised learning tasks Li & Wang (2019); Collins et al.
+(2021); Li et al.
+(2021b).
+These methods designed for supervised tasks focus on the heterogeneous problem in label distribution Diao et al.
+(2020) or feature distribution Li et al.
+(2022c), which cannot be directly adapted to the vision-language pre-training task.
+Recent works in representation-augmentation based uni-modal self-supervised federated learning Zhang et al.
+(2023a); Huang et al.
+(2022); Li et al.
+(2022b) employ additional communications on representations of local data to increase sample diversity.
+Such methods fail to protect data privacy, which is a vital concern in medical applications.
+Yan et al.
+(2023) utilizes image-side-only techniques to handle heterogeneity and thus may harm the learning of cross-modal alignment when adapted to the multi-modal setting.
+Zhuang et al.
+(2022); Li et al.
+(2021b) considers distillation-based methods yet ignores the direct modeling of cross-modal alignment.
+While the learned models can be biased and over-reliance on spurious correlations Saab et al.
+(2022) due to the data heterogeneity in the medical domain Krishnan et al.
+(2022), distributionally robust optimization Deng et al.
+(2020) can alleviate these issues by optimizing the group-wise worst-case performance on given objective Liu et al.
+(2022), thus can be flexibly adapted to various of federated learning tasks Han et al.
+(2023); Zhang et al.
+(2022a); Capitani et al.
+(2024).
 3 Problem Formulation
-We aim at mitigating the data heterogeneity issues in federally pre-training biomedical vision-language models to better capture the semantic alignment between the two modalities, with heterogeneous private datasets from different clients.
-In this section, we will formulate our problem set-up and assumptions.
-Please see Appendix A for the detailed preliminary.
-Federated Vision-Language Pre-training.
-In this paper, we consider the federated vision-language pre-training (VLP) problem in the biomedical domain.
-Suppose we are given $N$ clients $\{C_{i}\}_{i=1}^{N}$, each client has a corresponding local yet private dataset $\{D_{i}\}_{i=1}^{N}$ that can’t be easily shared due to privacy issues.
-Without access to centralized pre-training, federated pre-training aims to conduct vision-language representation learning across local clients, while still capture a robust and generalizable cross-modal alignment, by efficiently aggregating client models $\{M_{i}\}_{i=1}^{N}$ trained on their private datasets.
-Similar to conventional federated learning scenarios, our federated pre-training consists of multiple communication turns.
-For communication turn $t$, client models $\{M_{i}\}_{i=1}^{N}$ completed the same number of local iteration steps respectively, and the server receives and aggregates the weights sent from local models to produce the aggregated server model $M^{t}$.
-Then it will be sent back to clients and re-trained on $\{D_{i}\}_{i=1}^{N}$ separately to get new local models $\{M_{i}^{t+1}\}_{i=1}^{N}$.
-The federated vision-language pre-training is a more practical and general task compared to related works.
-Table 1 lists the most similar settings in previous works, and the difference between our task and theirs.
-Table 1: Comparison between related multi-modal federated learning tasks and our federated biomedical vision-language pre-training task.
-We organize the related works according to our key areas of interest.
+We aim at mitigating the data heterogeneity issues in federally pre-training medical vision-language models from heterogeneous private client datasets.
+In this section, we will formulate the set-up of the Federated Vision-Language Pre-training task, which aims to capture generalizable cross-modal alignment by leveraging diverse semantics from local datasets.
+Then we will introduce our assumption on multi-modal pre-training data and heterogeneity.
+Please see Appendix F for detailed preliminary.
+Federated Vision-Language Pre-training (FedVLP).
+We aim to federally pre-train medical vision-language models.
+Specifically, given $N$ clients $\{C_{i}\}_{i=1}^{N}$, FedVLP decentralizedly conducts vision-language representation learning across clients, where local models are learned only on the local client dataset.
+The goal is to aggregate the locally trained models on a central server to obtain a generalizable and unbiased global pre-trained model.
+Here, we assume each client has a local yet private dataset $\{D_{i}\}_{i=1}^{N}$ that can’t be easily shared due to privacy or other issues.
+During FedVLP, different from supervised federated learning, we focus on the transferability of the aggregated model.
+Different from supervised federated learning which is usually designed for a specific task, federated vision-language pre-training focuses on learning a generalizable and unbiased representation space, which can be transferred to various downstream tasks as mentioned in Sec 5.1.
+Employing an alternative updating approach, federated pre-training also consists of multiple communication turns between clients and the server , which is the same as conventional federated learning (FL) settings.
+For each communication turn $t$, client models $\{M_{i}^{t}\}_{i=1}^{N}$ completed local pre-training respectively and send local model weights to server, where all the local model weights are aggregated to obtain a server model $M^{t}$.
+$M^{t}$ will be broadcast to clients for updating local models in each communication turn, which completes the main part of the communication turn.
+In some FL methods, additional local computation and information transmissions may be conducted after clients have received the aggregated model, which will also be counted as a part of the communication.
+Table 18 in appendix lists the most similar settings, and their corresponding differences with ours.
+Formulation of Multi-Model Pre-training Data and Heterogeneity.
+We consider the multimodal datasets including vision and language modalities $X,Y$, consisting of image and text data samples respectively, $X=\{x_{i}\}_{i=1}^{n_{1}},Y=\{y_{i}\}_{i=1}^{n_{2}},n_{1},n_{2}\in N^{+}$.
+We assume there exist $m$-way discrete multi-label latent variables $U\in R^{m}$ that embed semantics shared across two modalities, and form multiple latent classes $\{u_{i}\}_{i=1}^{k},k\in N^{+}$.
+For example, disease labels of a given X-ray radiology-report pair, are latent semantic variables that connect these two modalities, which determines the pathology region of the radiology image and corresponding description in the diagnosis report.
+To model heterogeneity, we assume the distribution of latent variables is different across clients.
+Cross-modal alignment.
+We model cross-modal alignment as a module that outputs embeddings with high semantic similarity inspired by Arora et al.
 (2019).
-Heterogeneous clients.
+Therefore, we formulate the goal of the cross-modal alignment function, which is to minimize the distance between the distributions induced by vision and language representations.
+This could be achieved by employing a contrastive-based self-supervised learning object Huang et al.
+(2021b); Arora et al.
 (2019).
-Multi-modal clients.
-(3).
-Transferability.
-(4).
-Prior model-free.
-(5).
-No communication on public data or feature representations.
-Task hetero. m.m.c. transf. no prior no feat. com.
-(Yu et al., 2023a) $\checkmark$ $\checkmark$
-(Lu et al., 2023b) $\checkmark$ $\checkmark$ $\checkmark$
-(Chen et al., 2023b) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-(Lu et al., 2023a) $\checkmark$ $\checkmark$ $\checkmark$
-FedRGB $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Data Formulation.
-We consider the multimodal datasets including vision and language modalities $X,Y$, consisting of respective data samples $X=\{x_{i}\}_{i=1}^{n},Y=\{y_{i}\}_{i=1}^{m},n,m\in N^{+}$.
-We assume there exists latent variables $U$ that embed semantics across two modalities.
-For example, the disease category of a given X-ray radiology-report pair is a latent variable connecting these two modalities, which determines the pathology region of the radiology image and semantics in the diagnosis report.
-For heterogeneous local datasets, we assume the distribution of latent variables $P_{i}(U)$ is different across clients.
+Detailed analysis of this modeling can be seen in Appendix G.1.
 4 Method
-Figure 2: (a) Our proposed FedRGB upgrades backbone multi-modal pre-training strategies via introducing a robust guidance to local pre-training.
-This approach can reduce the distortion on feature encoders and learn better cross-modal alignment.
-(b) We employ a distribution-based min-max optimization to learn robust cross-modal alignment, which will be used as guidance in local pre-training.
-We propose a de-centralized framework to mitigate the data heterogeneity in federated biomedical vision-language pre-training, for learning an efficient vision-language model that can robustly capture reliable cross-modal semantic alignments.
-In this section, we first introduce our approach with a guidance-based local training strategy to reduce the distortion on encoders caused by fitting client datasets.
-Secondly, we will illustrate how our approach employ a distribution-based min-max optimization to learn unbiased cross-modal alignment.
-Algorithm 1 shows the pipeline of our method.
-4.1 Guidance-based Local Client Training
+Figure 2: (1) Federated pre-training strategies with uni-modality in previous works Zhuang et al.
+(2022); Jiang et al.
+(2023) often consider the representation or performance discrepancy, adapting them directly to multi-modal pre-training problem underestimates the necessity of modeling multi-modal interactions.
+(2) Our method considers the cross-modal alignment and binds local alignment to an unbiased one during local training.
+(3) In each communication turn, after local training and aggregating encoders, We employ a distribution-based min-max optimization to learn robust cross-modal alignment (3.b, 3.c).
+It will be used to guide local pre-training (c.1).
+To federally leverage heterogeneous local datasets to pre-train an efficient vision-language model, we propose a decentralized framework that can robustly capture reliable cross-modal semantic alignments across different clients.
+As shown in Fig.
+2, we employ the teacher alignment module to guide the local training, to learn factual diverse correlations between two modalities.
+In this section, we first introduce our approach to reduce distortion on global-aggregated features during fitting local dataset via binding cross-modal alignments, through a guidance-based local training strategy.
+Secondly, we will illustrate our approach of employing the distributionally robust optimization to learn unbiased cross-modal alignment.
+Details can be seen in Algorithm 1.
+4.1 Cross-Modal Alignment Binding via Guidance-Based Local Training
 As shown in our empirical findings (Sec.
-5.2), biomedical vision-language models pre-trained on heterogeneous client datasets are affected by the local correlations of the two modalities, which can’t be addressed by simply averaging the weights of local models.
-Furthermore, this mainly falls on the deep layers which act as alignment modules, where the heterogeneous alignment module would distort the learned encoder during local iterations as shown in Sec.
-7.
-Inspired by these findings, we introduce a teacher aligner to guide the representation space learned by local models to converge to a generalizable and robust one.
-Teacher alignment module.
-To alleviate the distortion and bias when learning local features encoders and alignment modules, we propose to utilize a teacher alignment module to generate anchor representations conditioned on the output feature vector of the encoder to encourage local models to learn a homogeneous representation space across clients.
-Suppose after one client-server communication turn, given the local dataset of client $c\in\mathcal{C}$ sampled from $(x^{c},y^{c})\sim D_{c}$, and teacher alignment module $g_{v}^{*},g_{l}^{*}$ corresponding to vision and language modality, we target at updating the feature encoders of image and text $f_{v},f_{l}$ and the corresponding alignment modules $g_{img},g_{txt}$, by training on the local client dataset $D_{c}$, under the guidance of teacher alignment modules.
-Guiding local training with robust aligner.
+5.2), medical vision-language models pre-trained on heterogeneous client datasets are affected by the local correlations of the two modalities, which can’t be addressed by simply averaging the weights of local models.
+Furthermore, this mainly falls on the deep layers which act as the cross-modal alignment module.
+As shown in Sec.
+8, we observe that a non-ideal alignment module would amplify bias in local datasets and enforce the encoder to learn biased features during local iterations and distort the global-aggregated features.
+Inspired by these findings, we design to separately handle the deep layers and shallow layers of the modality-specific network, which are considered to be the encoder and the alignment module, respectively.
+To encourage models to capture semantics in local datasets during local training, we keep updating the whole model during local training.
+To avoid feature distortion, we focus on the alignment module and bind the local cross-modal alignment to an ideal one.
+Binding cross-modal alignments on the representation space.
+Binding local cross-modal alignments can be straightforwardly solved by replacing the local alignment module with a frozen unbiased one.
+However, this is not practical due to our goal of encouraging models to learn semantics from local datasets.
+Besides, the unbiased alignment module of the current communication turn should also be updated.
+To enable cross-modal alignment binding with fully learnable local models, we employ an alternative approach to minimize the distance between the representation space projected by local alignment modules and an anchor representation space.
 As shown in Fig.
-2, given the image-text pairs from local dataset as input $(x^{c},y^{c})\sim D_{c}$, we denote the feature vectors output from the feature encoder modules as $(v^{c},l^{c})$.
-Then getting through the alignment module, the aligned representations output from local aligners are denoted by $(z_{I}^{c},z_{T}^{c})$, while those from teacher aligners are $(z^{c*}_{I},z^{c*}_{T})$.
-The goal is to reduce the discrepancy between two sets of aligned representations produced by local and teacher aligners, so that we can keep locally trained model will not be misled by the bias from local datasets and all local clients can converge to learn a similar embedding space that captures globally reliable and generalizable cross-domain alignment.
-The teacher guidance loss can be formulated as below:
-$\footnotesize\mathcal{L}_{guide}=\sum_{i=1}^{N_{c}}(z^{c*}_{I,i}-z^{c}_{I,i})^{2}+\sum_{i=1}^{N_{c}}(z^{c*}_{T,i}-z^{c}_{T,i})^{2}$ (1)
-, where $N_{c}$ refers to the number of samples in client the $c$.
-We combine the guidance loss with the optimization goal $\mathcal{L}_{task}$ of the initial biomedical vision-language pre-training method, which includes e.g.
-image-text contrastive loss in ConVIRT (Zhang et al., 2022b).
-The overall loss function for local client training can be written as:
-$\footnotesize\mathcal{L}_{local}=\mathcal{L}_{guide}+\alpha\mathcal{L}_{task}$ (2)
-where $\alpha$ is the hyper-parameter that balances between guidance regularization and training task in local client updating.
-4.2 Learning Unbiased Cross-modal Alignment via Distribution-based Optimization
-In order to guide the local client training, we need to provide a teacher alignment module as mentioned in the last section, which can capture globally unbiased and generalizable cross-domain alignment.
+2, given input image-text pairs from dataset of client $c\in\mathcal{C}$ as $(x^{c},y^{c})\sim D_{c}$, where c is the index of client, the feature vectors output from the image and text encoder modules can be denoted as $(I^{c},T^{c})$, respectively.
+$(I^{c},T^{c})$ will be aligned by alignment modules, and the aligned representations output from local aligners and anchor representations output from unbiased aligner are denoted by $(z_{I}^{c},z_{T}^{c})$, $(z^{c*}_{I},z^{c*}_{T})$ respectively.
+During the local training, given the learning object of aligning two modalities $L_{task}(x,y)$ (e.g.
+image-text contrastive loss in ConVIRT Zhang et al.
+(2022b)), we have:
+Remark 1
+Consider $L$-lipschitz learning object $L_{\text{task}}:\mathbb{R}^{2d}\to\mathbb{R},L_{\text{task}}>0,d\in\mathbb{N}^{+}$.
+Minimizing $L_{\text{task}}(z^{c*}_{I},z^{c*}_{T})$ can be reformulated as: $\min_{\theta}\left(L_{\text{task}}(z_{I}^{c},z_{T}^{c})+L\left(\|z_{I}^{c}-z^{c*}_{I}\|_{2}^{2}+\|z_{T}^{c}-z^{c*}_{T}\|_{2}^{2}\right)\right)$, where $\theta$ represents the model parameters.
+See detailed analysis in Appendix G.1.
+Guiding local training with robust alignment module.
+To bind local cross-modal alignments, motivated by Remark 1, we can introduce a robust teacher alignment module to generate anchor representations conditioned on the output feature vector of the encoder, and leverage them for guiding local training.
+Specifically, given the local dataset of client $c\in\mathcal{C}$ sampled from $(x^{c},y^{c})\sim D_{c}$, and teacher alignment module $g_{I}^{*},g_{T}^{*}$, we target at updating the feature encoders $f_{I},f_{T}$ and the corresponding alignment modules $g_{I},g_{T}$, by training on the local client dataset $D_{c}$, under the guidance of teacher alignment modules.
+Following the above analysis, we aim to reduce the discrepancy between two sets of aligned representations produced by local and teacher alignment modules.
+Thus, given number $N_{c}$ of samples in client $c$, the teacher guidance loss can be formulated as below:
+$\scriptsize\mathcal{L}_{guide}=\sum_{i=1}^{N_{c}}(z^{*}_{I,i}-z_{I,i})^{2}+\sum_{i=1}^{N_{c}}(z^{*}_{T,i}-z_{T,i})^{2}.$ (1)
+We combine the guidance loss with the optimization goal $\mathcal{L}_{task}$ of the initial medical vision-language pre-training method, which includes e.g.
+image-text contrastive loss in ConVIRT Zhang et al.
+(2022b).
+The overall loss function for guidance-based local client training can be written as:
+$\scriptsize\mathcal{L}_{local}=\mathcal{L}_{guide}+\alpha\mathcal{L}_{task}$ (2)
+where $\alpha$ is the hyper-parameter that balances between guidance regularization and training tasks.
+4.2 Learning Unbiased Cross-modal Alignment via Distributionally Robust Optimization
+As mentioned in section 4.1, we aim to utilize an anchor space to guide the local client training.
+To achieve this, we need to obtain a teacher alignment module that can capture globally unbiased and generalizable cross-domain alignment.
 How can we obtain such a teacher guidance model?
-We propose to learn the unbiased cross-domain alignment by considering the robustness across various distributions of local clients.
-Detailed problem formulation and description could be seen in Appendix B.
-A teacher alignment module should avoid distorting the aggregated feature encoder while learning the unbiased cross-modal alignments.
-Additionally, we introduce a much harder task to force the module to be more robust to biased alignment, by considering the influence of the biased alignment module on the feature encoder.
-Furthermore, to learn a teacher model that performs well on varies data distributions, we minimize distortion degree under the most biased local alignment with the “worst-case” local data distribution.
+We propose to learn such unbiased cross-modal alignment by considering the robustness across various distributions of local clients.
+Specifically, the teacher alignment module should encourage the model to learn well-aligned latent space while resulting in less feature distortion on the aggregated feature encoder during fitting local datasets.
+Furthermore, the robustness requires the teacher model to perform well on various data distributions.
+These insights motivate us to learn an alignment module that can well align representations and reduce feature discrepancy under various local clients, which would even well tackle the most severe distortion during local training.
+This motivation inspires us to model the problem as a worst-case optimization problem, which can be solved by robust optimization (DRO) Cotter et al.
+(2019); Namkoong & Duchi (2016).
+Detailed problem formulation and description can be seen in Appendix F.
 Learning robust cross-modal alignment through distributionally robust optimization.
-Inspired by the distributionally robust optimization (DRO) (Cotter et al., 2019), we model the de-centralized pre-training framework as a two-stage DRO learning process.
-In the first stage, the local models are trained with $\mathcal{L}_{local}$ (Eq.
-2), while the whole model is jointly optimized with the guidance of the teacher alignment module.
-In the second stage, we employ the decentralized DRO approach to learn a robust teacher alignment module that performs well conditioned on the worst-case local models where the feature encoder is distorted the most severely.
-Learning object of alleviating distortion in local triaing.
-After finishing the local training in each communication round, we aim to employ an extra de-centralized learning process in the client-side to optimize the alignment module on the worst-case distribution of an uncertainty set constructed by the local aligner and local data distribution, denoted as $P_{i}$ for simplicity.
-The uncertainty set can be written as:
-$\footnotesize Q^{c}:\{D(Q||P^{c})\leq\rho\},P^{c}=\sum_{i=1}^{N}w_{i}P_{i},w\in\Delta_{N-1}$ (3)
-where $N$ is the number of clients, $\Delta$ is a $(N-1)$-dim simplex, $D$ is a discrepancy measure, and $\rho$ is the uncertainty radius that measures the uncertainty degree of the unseen distributions and influences model’s generalization ability.
-The details of the $P_{i}$ can be seen in Appendix B.
-With the teacher alignment module denoted as $g^{*}$ and the aggregated encoder $f^{*}$ after the first stage training, the optimization goal can be constructed by distributions $P_{i}$ of the support set:
-$\footnotesize min_{g^{*}}\ max_{\textbf{w}\in\Delta_{N-1}}\sum_{i=1}^{N}w_{i}\ \mathcal{L}_{dro}(g^{*},P_{i},f^{*}),s.t.D(N\textbf{w}||\textbf{1})\leq\rho$ (4)
-,where $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$ are the losses calculated on the local clients respectively.
-While we learn $g^{*}$ by gradient descent, employing an alternative optimization approach, the mirror gradient descent process of optimizing w could be solved efficiently (Zhang et al., 2023b; Duchi et al., 2008):
-$\footnotesize{w_{i}^{(r+1)}}=\text{Proj}_{\mathcal{Q}{c}}\left({w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}}/\sum_{i=1}^{N}{w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}},\rho\right)$ (5)
-where $\text{Proj}_{\mathcal{Q}{c}}$ refers to the projection to $\{w:D(N\textbf{w}||\textbf{1})\leq\rho\}$, $v_{i}$ represents the loss $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$, $r$ refers to the communication turn.
-Finally, we formulate $\mathcal{L}_{dro}$ to measure the performance of the teacher alignment module of alleviating distortion on the encoder.
-Based on Eqs 3 and 4, we update the teacher alignment module in local clients with weights w.
-Specifically, given a client $c$, we use $f^{*}$ learned in the first stage of the current communication turn to generate features $(\tilde{v^{c}},\tilde{l^{c}})$ which are not distorted by the local aligner during the second stage’s training.
-And a learnable copy of $f^{*}$ to generate distorted features $(v^{c},l^{c})$.
-The distorted features are input to the frozen local aligner $g$ to calculate the task loss $L_{task}$, and are also input to the teacher aligner $g^{*}$ to get calibrated representations $(z_{I}^{c*},z_{T}^{c*})$.
-Meanwhile, $(\tilde{v^{c}},\tilde{l^{c}})$ are input to $g^{*}$ to get representation $\tilde{(z_{I}^{c*}},\tilde{z_{T}^{c*}})$ as anchors for $(z_{I}^{c*},z_{T}^{c*})$ projected by teacher aligner using distorted features, for regularization.
-The learning object of the second stage could be written as:
-$\footnotesize\mathcal{L}_{dro}=\beta(\sum_{i=1}^{N_{c}}(\tilde{z_{I,i}^{c*}}-z_{I,i}^{c*})^{2}+\sum_{i=1}^{N_{c}}(\tilde{z_{T,i}^{c*}}-z_{T,i}^{c*})^{2})+\mathcal{L}_{CL}+\mathcal{L}_{task}$ (6)
-,where $\mathcal{L}_{CL}$ is the contrastive loss (Radford et al., 2021) between $\tilde{(z_{I}^{c*}},\tilde{z_{T}^{c*}})$, $\mathcal{L}_{task}$ refers to the loss function of the backbone pre-training method and $N_{c}$ refers to the number of samples in the client $c$.
+Following the analysis above, we model the de-centralized pre-training framework as a two-stage DRO learning process.
+In the first stage, the whole model is jointly optimized with $\mathcal{L}_{local}$ (Eq.
+2), and the aggregated encoder $f^{*}$ is obtained by server averaging.
+In the second stage, we employ the decentralized DRO to learn a robust teacher alignment module $g^{*}$ that performs well conditioned on the worst-case local models and local data distribution where the feature encoder is distorted the most severely.
+We employ an extra de-centralized learning process on the client side to optimize the alignment module on the worst-case distribution of an uncertainty set Namkoong & Duchi (2016) constructed by the local clients, with the distributions denoted as $P_{i},i\in[1,...,N_{c}]$ for simplicity.
+The details of the uncertainty set can be seen in the Appendix Sec.
+F.
+Given teacher alignment module $g^{*}$ and the aggregated encoder $f^{*}$, the optimization goal can be constructed based on the support distribution $P_{i}$ in the uncertainty set:
+$\scriptsize\min_{g^{*}}\ \max_{\textbf{w}\in\Delta_{N-1}}\sum_{i=1}^{N}w_{i}\ \mathcal{L}_{dro}(g^{*},P_{i},f^{*}),s.t.D(N\textbf{w}||\textbf{1})\leq\rho$ (3)
+,where $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$ is calculated on local clients. We will illustrate it in the following part.
+Optimizing teacher alignment module via alternating min-max game.
+To match the newly learned semantics from local datasets during local training of each communication turn $r$, we aim to continually update the teacher alignment module.
+For each turn $r$, we employ an alternative optimization approach.
+We firstly complete the local training, and optimize w via the mirror gradient descent process that could be solved efficiently Namkoong & Duchi (2016); Zhang et al.
+(2023b); Duchi et al.
+(2008):
+$\scriptsize{w_{i}^{(r+1)}}=\text{Proj}_{\mathcal{Q}{c}}\left({w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}}/\sum_{i=1}^{N}{w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}},\rho\right)$ (4)
+where $\text{Proj}_{\mathcal{Q}{c}}$ refers to the projection to $\{w:D(N\textbf{w}||\textbf{1})\leq\rho\}$, $v_{i}$ represents the loss $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$.
+The weights w will be used in updating $g^{*}$ following Eq.
+3 in local clients.
+Finally, we propose the $\mathcal{L}_{dro}$ to learn a $g^{*}$ that can well align representations of the two modalities and reduce the discrepancy between anchor and local representations projected by $g^{*}$ and local alignment module $g$ during local training.
+To reduce the discrepancy, we directly minimize the distance of anchor and local representations modality-wise.
+Specifically, for client $c$, we use $f^{*}$ to generate features $(\tilde{I},\tilde{T})$ which are not distorted by the local aligner during the second stage’s training, and a learnable copy to generate distorted features $(I,T)$.
+$(\tilde{I},\tilde{T})$ and $(I,T)$ are input to $g^{*}$ to get $(z_{I}^{*},z_{T}^{*})$ and $\tilde{(z_{I}^{*}},\tilde{z_{T}^{*}})$ respectively.
+Then we minimize the L2 distance of them.
+To guarantee the alignment in latent space, we utilize multi-modal contrastive loss Radford et al.
+(2021); Zhang et al.
+(2022b) $\mathcal{L}_{CL}$ between $\tilde{(z_{I}^{*}},\tilde{z_{T}^{*}})$ to guarantee to learn well-aligned latent space.
+To simulate the distortion by local training, we input $(I,T)$ into $g$ to get $(z_{I},z_{T})$ to calculate $\mathcal{L}_{task}$.
+The learning object of the second stage of client $c$ could be written as:
+$\scriptsize\mathcal{L}_{dro}=\beta(\sum_{i=1}^{N_{c}}(\tilde{z_{I,i}^{*}}-z_{I,i}^{*})^{2}+\sum_{i=1}^{N_{c}}(\tilde{z_{T,i}^{*}}-z_{T,i}^{*})^{2})+\mathcal{L}_{CL}+\mathcal{L}_{task}.$ (5)
 5 Experiment
 5.1 Experiment Setting
-5.1.1 Pre-training setup
-Following (Wang et al., 2022), we utilize the MIMIC-CXR (Bigolin Lanfredi et al., 2022) dataset for biomedical vision-language pre-training.
+Pre-training setup.
+Following Wang et al.
+(2022), we utilize the MIMIC-CXR Bigolin Lanfredi et al.
+(2022) dataset for medical vision-language pre-training.
 During the de-centralized pre-training process, local clients only have access to their private datasets.
 Following Yan et al.
-(2023), we employ the Latent Dirichlet Allocation (LDA) (Blei et al., 2003) to divide the MIMIC-CXR dataset based on a 5-way disease attribute to construct the heterogeneous client datasets.
+(2023), we employ the Latent Dirichlet Allocation (LDA) Blei et al.
+(2003) to divide the MIMIC-CXR dataset based on a 5-way disease attribute to construct the heterogeneous client datasets.
 We set the heterogeneity degree in the LDA algorithm to be 1.
-The client datasets are highly heterogeneous as shown in Figure 3 and the appendix.
+The client datasets are highly heterogeneous as shown in Fig.
+4 and the Appendix.
 We divide the MIMIC-CXR into 5 heterogeneous subgroups to construct 5 clients.
-Each divided dataset consists of train splits and validation splits based on the notation of the MIMIC-CXR.
+Each divided dataset consists of train splits and test splits based on the notation of the MIMIC-CXR.
 For the pre-training reported in the main table, we set the number of communications to 25, and we randomly sampled 50 batches of data from the client datasets for each communication.
-For more details, please see the appendix.
-For analysis experiments on out-of-distribution generalization ability, we utilize the EHRXQA (Bae et al., 2023) to evaluate the transferability of the pre-trained model to the unseen domain.
-Figure 3: The descriptive statistic of the divided 5 subgroups, it can be seen that they are highly heterogeneous.
-5.1.2 Downstream tasks
-We conduct several downstream tasks following (Wang et al., 2022) to evaluate the generalization ability of the pre-trained model.
-For more details, please see the appendix.
-Few-shot classification.We test their performance on multiple image classification benchmarks: (1) RSNA Pneumonia Detection (RSNA) (Shih et al., 2019).
-(2) Covidx (Wang et al., 2020).
+For more details, please see the Appendix B.1.
+Downstream tasks.
+We conduct several downstream tasks following Wang et al.
+(2022) to evaluate the generalization ability of the pre-trained model.
+For more details, please see the Appendix B.2.
+(1).
+Few-shot classification.
+We test their performance on multiple image classification benchmarks: (a).
+RSNA Pneumonia Detection (RSNA) Shih et al.
+(2019).
+(b) Covidx Wang et al.
+(2020).
 We fine-tune our pre-trained model with an additional linear layer on 1%, 10% percentage of the training dataset, and evaluate the classification accuracy.
+(2).
 Medical image segmentation.
-We conduct experiments following (Wang et al., 2022) on medical image segmentation on the RSNA (Wang et al., 2020) benchmark.
+We conduct experiments following Wang et al.
+(2022) on medical image segmentation on the RSNA Wang et al.
+(2020) benchmark.
 We freeze the encoder and fine-tune a U-Net decoder using 1%, 10% of the training data, and then use the Dice score for evaluation.
+(3).
 Image retrieval.
@@
 We utilize the top-1 recall accuracy and top-5 recall accuracy to evaluate the pre-trained models.
-5.1.3 Backbones and Baselines
-We focus on upgrading biomedical vision-language pre-training (VLP) strategies to heterogeneous federated learning scenes.
-Since federated heterogeneous VLP is a new topic, we adopt existing biomedical VLP baselines to federated strategies (Lu et al., 2023a).
-Specifically, we consider adopting simple language-image contrastive alignment (ConVIRT) (Zhang et al., 2022b; Radford et al., 2021), global-local language-image contrastive alignment(GLoRIA) (Huang et al., 2021), and Multi-Granularity Cross-modal Alignment (MGCA) (Wang et al., 2022).
-Since the federated multi-modal per-training is a novel setting, we adapted FedMAE (Yan et al., 2023), FedEMA (Zhuang et al., 2022), and FedX (Han et al., 2022) for comparison.
+Backbones and Baselines.
+We focus on enabling medical vision-language pre-training (VLP) methods to be applied to heterogeneous federated learning scenes and have considered the generalization ability of FedAID on different backbone VLP methods.
+Since federated heterogeneous VLP is a new topic, we adopt existing medical VLP baselines to federated strategies Lu et al.
+(2023a).
+Specifically, we consider adopting contrastive learning based methods: simple language-image contrastive alignment (ConVIRT) Zhang et al.
+(2022b); Radford et al.
+(2021), global-local language-image contrastive alignment(GLoRIA) Huang et al.
+(2021a), and Multi-Granularity Cross-modal Alignment (MGCA) Wang et al.
+(2022).
+Since the federated multi-modal per-training is a novel setting, we adapted FedMAE Yan et al.
+(2023), FedEMA Zhuang et al.
+(2022), FedMOON Li et al.
+(2021b) for comparison.
 These are self-supervised learning methods in the federated learning domain which also focus on tackling the data heterogeneity.
-For basic federated learning baselines, we consider simple averaging (FedAvg) (McMahan et al., 2017; Lu et al., 2023a), centralized training (Local), and decentralized training (Global).
-For fair comparisons, we combine each federated learning strategy with different biomedical VLP methods.
+For basic federated learning baselines, we consider simple averaging (FedAvg) McMahan et al.
+(2017); Lu et al.
+(2023a), decentralized training (Local), and centralized training (Global).
+For fair comparisons, we combine each federated learning strategy with different medical VLP methods.
 For baselines pre-trained in Local strategy, we report the averaged performance of the local models.
@@
 To enable uni-modal self-supervised learning baselines to be applied to our setting, we employ the image-text contrastive loss with the same hyperparameters to our method for fairness.
-We utilize a ViT-base (Dosovitskiy et al., 2020) as our vision encoder and utilize the Bert-base (Devlin et al., 2018) as our text encoder.
-We pre-process the input following (Wang et al., 2022).
+We utilize a ViT-base Dosovitskiy et al.
+(2020) as our vision encoder and utilize the Bert-base Devlin et al.
+(2018) as our text encoder.
+We pre-process the input following Wang et al.
+(2022).
 We utilize an additional transformer layer from ViT-base and Bert-base as the align module for vision and language respectively.
 5.2 Empirical Finding
-In this section, we will demonstrate our key empirical findings for federated multi-modal pre-training under heterogeneous client datasets, which actually motivates us to propose our Federated distributional Robust Guidance-Based (FedRGB).
+In this section, we will demonstrate our key empirical findings for federated multi-modal pre-training under heterogeneous client datasets, which actually motivates us to propose our Federated distributional Robust Guidance-Based (FedAID).
+We conduct experiments on the image-text retrieval task, which reflects the ability to capture vision-language alignments.
+Due to the page limitation, here we list our key findings, detailed results can be seen in the Appendix A:
+  • FedAvg can aggregate semantics from local clients, but the performance would be damaged by data heterogeneity as shown in Fig. 5.
+  • Fig. 6 shows retraining server model on local datasets would distort its feature encoders.
+  • Fig. 7 shows diverse info. is aggregated in shallow layers that extract feature semantics.
+  • De-centralized pre-trained deep layers that act as the align module is harmful for extracting diverse semantics as shown in Fig. 8.
+Overall, from empirical findings, we conclude that federated multimodal pre-training is sensitive to data heterogeneity which results in learning distorted cross-modal alignment.
+Simply averaging local model weights does not solve this issue essentially.
+Thus, we are motivated to propose our method.
+Table 1: Downstream task performance. We report the few-shot classification accuracy on Covid and RSNA, the retrieval accuracy on Chest imaGenome, and the Dice score for segmentation on RSNA.
+Strategy RSNA (cls.) Covid (cls.) RSNA (seg.) Chest imaG. (ret.)
+$1\%$ $10\%$ $1\%$ $10\%$ $1\%$ $10\%$ Rec.@1 Rec.@5 Wst.@1 Wst.@5
+ImageNet Init ViT 79.02 80.62 63.75 83.75 60.09 67.02 3.64 16.46 2.86 14.17
+FedMAE 80.81 81.06 77.50 85.75 69.48 73.38 8.29 33.03 6.92 30.49
+FedEMA 82.81 83.19 79.25 86.50 70.94 73.65 24.01 67.05 21.85 62.44
+FedMOON 82.51 83.21 77.75 89.25 68.99 71.28 27.85 70.88 25.34 67,19
+FedAvg 83.19 83.31 78.00 88.50 69.59 71.50 28.83 72.12 25.34 66.68
+FedAID (Ours) 83.21 83.54 82.50 90.00 71.51 74.08 29.49 73.03 26.99 67.65
+Centralized 83.39 84.61 82.50 92.00 72.60 76.35 41.45 84.22 38.55 79.98
+Table 2: Downstream task performance on different multi-modal pre-training backbone methods.
+Strategy Method RSNA (cls.) Covid (cls.) RSNA (seg.) Chest imaG. (ret.)
+$1\%$ $10\%$ $1\%$ $10\%$ $1\%$ $10\%$ Rec.@1 Rec.@5 Wst.@1 Wst.@5
+Decentralized GLoRIA 82.34 82.89 77.90 88.75 71.07 72.13 17.15 52.50 15.19 48.73
+FedAvg GLoRIA 83.16 83.28 77.50 89.00 71.37 72.38 29.90 73.80 27.78 69.46
+FedAID GLoRIA 83.39 83.51 78.50 89.25 72.43 73.50 30.94 73.89 27.86 69.53
+Centralized GLoRIA 84.01 84.74 82.25 91.75 73.63 73.64 41.69 84.04 39.00 80.52
+Decentralized MGCA 81.90 82.71 77.80 87.55 62.77 70.20 15.21 50.44 13.37 45.39
+FedAvg MGCA 82.63 83.54 75.75 88.25 70.06 71.39 29.33 73.71 26.81 70.49
+FedAID MGCA 82.76 83.61 77.75 88.25 70.71 72.52 29.96 74.15 27.11 69.53
+Centralized MGCA 83.96 84.49 79.50 89.50 70.71 72.52 39.89 83.50 36.88 80.28
+5.3 Main Result
+Our method learns robust and enriched cross-modal alignment and has better transferability.
+Table 1 shows the image-text retrieval results.
+Here we utilize the ConVIRT as the backbone pre-training method.
+Both average and worst-client accuracies of our method are higher than baseline’s, which means our model can not only capture more cross-modal alignment but optimize the performance on the worst client dataset, which verifies the robustness and fairness of our DRO-based strategy.
+Table 1 shows the results of the classification and segmentation downstream tasks.
+Our method beats other baseline strategies on each task and can be comparable to the centralized trained model, which acts as the upper bound.
+Therefore, our method has better transferability.
+Previous single-modality methods cannot be adapted directly for multi-modal data.
+Table 1 shows the performance of adapted self-supervised federated learning methods which focus on single-modality.
+From the results, we observe that baselines have shown transferability on visual downstream tasks.
+However, in the multi-modal retrieval task, our method beats these baselines by a large margin, which indicates that previous single-modality methods cannot be easily adapted directly for multi-modal data.
+We conjecture this is because they have ignored modeling the cross-modal alignment.
+Furthermore, we think FedAvg is still a strong baseline compared to other adapted methods, as we observed in the experiments.
+That’s because FedAvg doesn’t hurt the ability to learn cross-modal alignment of backbone pre-training models by adding augmented losses.
+Our method can be transferred to multiple multi-modal pre-training methods.
+Table 2 shows the downstream task performance of the MGCA and the GLoRIA backbone pre-training method when combined with our strategy.
+Our method has successfully upgraded MGCA and GLoRIA to adapt to the heterogeneous federated multi-modal pre-training scene, as shown in the significant improvement of the classification and segmentation downstream tasks.
+5.4 Analysis experiment
+In this section, we conduct analysis experiments using ConVIRT as backbone pre-training method.
+Table 3: Ablation results. We report the few-shot classification accuracy and the retrieval accuracy.
+Alignment Module Guidance $1\%$ (cls.) Chest imaG. (ret.)
+Covid RSNA Rec.@1 Rec.@5 Wst.@1 Wst.@5
+Simple-Average L2 83.14 79.75 28.31 72.18 26.26 66.56
+DRO learned Contrastive 82.91 78.25 28.81 71.03 25.95 66.42
+DRO learned L2 83.21 80.50 29.49 26.99 73.03 67.65
+[…]
+Robust teacher align module is better than simply averaged one.
+To investigate the efficiency of the robust align module, we replace it with an align module that is simply averaged by local ones and pre-train a new model.
+As shown in table 3, the downstream performance of the replaced one is lower than the original one.
+This means the robust align module can enhance the ability to capture cross-modal alignments and learn transferable representation.
+Strong regularization is better than weak regularization for guiding the updation.
+To investigate the effect of the guidance from the robust align module, we replace the L-2 regularization with contrastive loss as a weaker regularization and pre-train a new model.
+As shown in table 3, the weaker guidance leads to lower classification accuracy and retrieval accuracy in downstream tasks.
+Guidance hyper-parameter adjusts the average and disparity of client performances.
+Fig.
+3.a shows that higher guidance achieves smaller disparity while bringing less harm on average performance.
+We changed the hyperparameter $\alpha$ from 2 to 10, the disparity of the retrieval accuracy on test sets of different clients decreased significantly while there was only a slight loss on the averaged performance.
+On the other hand, Fig.
+3.b shows that increasing $\alpha$ wouldn’t result in a penalty in the transferability of the pre-trained backbone, the downstream task performances are maintained at a high level as we enlarge the guidance.
+Detailed results can be seen in Table 13.
+Smaller radius of uncertain set brings higher generalization and transferability on downstream task.
+Fig.
+3.c shows the downstream performance of models pre-trained with different radius of uncertainty set in the DRO process.
+As the $\rho$ increases, the performance of fine-tuning-based downstream tasks shows an upward trend.
+Detailed results can be seen in Table 7.
+Convergence. We show the transferability of our models pre-trained with different communication turns in Fig. 3.d. As commun. turn reaches 20, the downstream performances tend to converge.
+We have also conducted experiments to explore other properties of our proposed method.
+We have found that (1).
+Our method has successfully avoided the distortion caused by heterogeneous datasets during the local training.
+(2).
+Bringing more diverse clients for the federated pre-training is likely to improve the performance of the pre-trained model.
+(3).
+Our method is more generalizable on out-of-distribution multi-modal X-Ray datasets.
+Detailed results can be seen in Appendix C
+6 Conclusion
+Data limitation is a long-standing problem in the multi-modal learning domain.
+Despite federated learning can leveraging datasets from multiple sources while guaranteeing privacy issues, its performance would be damaged by data heterogeneity.
+Inspired by our empirical findings on the impact of heterogeneity on federated multi-modal learning, we propose the FedAID framework to multigate heterogeneity for federated medical vision-language pre-training.
+The efficiency of our method has been verified by comprehensive experiments.
+While introducing representation from other clients might bring larger improvement, we still consider the most privacy-preserved setting where representations are not transmissible.
+Further work could explore how to introduce data diversity while keeping local data private.
+7 Acknowledgement
+Zitao Shuai and Liyue Shen acknowledges support from U-M MIDAS PODS Grant and U-M MICDE Catalyst Grant, and computing resource support from NSF ACCESS Program.
+This work used NCSA Delta GPU through allocation CIS230133 and ELE230011 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services &\&& Support (ACCESS) program, which is supported by National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296.
+Appendix A Empirical Finding
+In this section, we will demonstrate our key empirical findings for federated multi-modal pre-training under heterogeneous client datasets, which actually motivates us to propose our FedAID.
 We conduct experiments on the image-text retrieval task, which reflects the ability to capture vision-language alignments.
 We utilize the ConVIRT backbone strategy and consider the performance of the centralized pre-trained model (Global), the server model pre-trained in FedAvg(FedAvg), and the average performance of decentralized pre-trained local models (Local.avg.).
+Figure 4: The disease distribution of the divided 5 subgroups, they are highly heterogeneous.
 Figure 5: The comparison of retrieval acc. on each client denoted as $\{C_{i}\}_{i=1}^{5}$, of centralized, FedAvg, and averaged acc. of decentralized pre-trained models.
@@
 FedAvg can aggregate semantics from local clients, but the performance would be damaged by data heterogeneity.
-Figure 4 shows retrieval accuracies of the strategies we have considered, where the FedAvg pre-trained model achieves higher accuracies than those of decentralized ones with a large margin.
+Fig.
+5 shows retrieval accuracies of the strategies we have considered, where the FedAvg pre-trained model achieves higher accuracies than those of decentralized ones with a large margin.
 However, the centralized pre-trained model acts as an upper bound and there still exists a great space for improvement for FedAvg.
@@
 Retraining the server model on local datasets would distort its feature encoders.
-Figure 5 demonstrates the effect of distortion by fitting local cross-modal alignments of the corresponding client datasets.
+Fig.
+6 demonstrates the effect of distortion by fitting local cross-modal alignments of the corresponding client datasets.
 We consider a set of models that finished the 25th local iteration and their corresponding aggregated server model.
@@
 We re-train the first 4 shallow layers of the aggregated server model on local clients’ datasets for 500 iteration steps.
-As shown in figure 6, the performance of retrained models decreases a lot, which means shallow layers that extract feature semantics are distorted by heterogeneous local datasets.
+As shown in Fig.
+7, the performance of retrained models decreases a lot, which means shallow layers that extract feature semantics are distorted by heterogeneous local datasets.
 Table 15 in the appendix shows the detailed results.
@@
 We re-train the first 4 shallow layers of the decentralized pre-trained model on the union of client datasets for 500 steps.
-Although the performance has successfully increased, there is still a large gap between that of the FedAvg pre-trained baselines and them as shown in Figure 7.
+Although the performance has successfully increased, there is still a large gap between that of the FedAvg pre-trained baselines and them as shown in Fig.
+8.
 Therefore, the frozen deep layers prevent the model from learning more diverse semantics from the scaled dataset.
@@
 Simply averaging the local model weights does not solve this issue essentially.
-Thus, we are motivated to propose our FedRGB.
-Table 2: Downstream task performance. We report the few-shot classification accuracy on Covid and RSNA, the retrieval accuracy on Chest imaGenome, and the Dice score on RSNA.
-Strategy RSNA (cls.) Covid (cls.) RSNA (seg.) Chest imaG. (ret.)
-$1\%$ $10\%$ $1\%$ $10\%$ $1\%$ $10\%$ Rec.@1 Rec.@5 Wst.@1 Wst.@5
-Decentralized 82.27 82.83 79.45 88.25 67.44 70.75 15.51 51.12 13.62 45.97
-FedX        (Han et al., 2022) 80.16 81.17 80.50 89.25 71.27 71.61 3.22 16.76 1.58 11.94
-FedMAE (Yan et al., 2023) 76.97 80.66 42.25 49.00 69.48 73.38 5.05 23.91 4.34 20.86
-FedEMA (Zhuang et al., 2022) 82.81 83.19 79.25 86.50 70.94 73.65 24.01 67.05 21.85 62.44
-FedAvg   (Lu et al., 2023a) 83.19 83.31 78.00 88.50 69.59 71.50 28.83 72.12 25.34 66.68
-FedRGB (Ours) 83.21 83.54 82.50 90.00 71.51 74.08 29.49 73.03 26.99 67.65
-Centralized 83.39 84.61 82.50 92.00 72.60 76.35 41.45 84.22 38.55 79.98
-Table 3: Downstream task performance on different multi-modal pre-training backbone methods.
-Strategy Method RSNA (cls.) Covid (cls.) RSNA (seg.) Chest imaG. (ret.)
-$1\%$ $10\%$ $1\%$ $10\%$ $1\%$ $10\%$ Rec.@1 Rec.@5 Wst.@1 Wst.@5
-Decentralized GLoRIA 82.34 82.89 77.90 88.75 71.07 72.13 17.15 52.50 15.19 48.73
-FedAvg GLoRIA 83.16 83.28 77.50 89.00 71.37 72.38 29.90 73.80 27.78 69.46
-FedRGB (Ours) GLoRIA 83.39 83.51 78.50 89.25 72.43 73.50 30.94 73.89 27.86 69.53
-Centralized GLoRIA 84.01 84.74 82.25 91.75 73.63 73.64 41.69 84.04 39.00 80.52
-Decentralized MGCA 81.90 82.71 77.80 87.55 62.77 70.20 15.21 50.44 13.37 45.39
-FedAvg MGCA 82.63 83.54 75.75 88.25 70.06 71.39 29.33 73.71 26.81 70.49
-FedRGB (Ours) MGCA 82.76 83.61 77.75 88.25 70.71 72.52 29.96 74.15 27.11 69.53
-Centralized MGCA 83.96 84.49 79.50 89.50 70.71 72.52 39.89 83.50 36.88 80.28
-5.3 Main Result
-Our method learns robust and enriched cross-modal alignment and has better transferability.
-Table 2 shows the image-text retrieval results.
-Here we utilize the ConVIRT as the backbone pre-training method.
-We report the average recall accuracy and the worst-case recall accuracy.
-The two sets of acc.
-are higher than those of FedAvg, which means our model can not only capture more cross-modal alignment but optimize the performance on the worst client dataset, which verifies the robustness and fairness of our DRO-based strategy.
-Table 2 shows the results of the classification and segmentation downstream tasks.
-In general, our method has better transferability than the traditional FedAvg strategy, most of the compared metrics of our method are higher than those of the baseline, and are closer to those of the centralized pre-trained model, which is the upper bound baseline.
-Other self-supervised baselines that aim to tackle the data heterogeneity have weaker performance compared to FedAvg, we analyze that this is due to the lack of the modeling of heterogeneous multi-modal alignment.
-Previous single-modality methods cannot be adapted directly for multi-modal data.
-Table 2 shows the performance of adapted self-supervised federated learning methods which focus on single-modality.
-From the results, we observe that our method beats these baselines by a large margin, which indicates that previous single-modality methods cannot be easily adapted directly for multi-modal data.
-Particularly, FedMAE and the FedX perform much worse than other federated learning strategies.
-We conjecture this is because both of them have ignored modeling the cross-modal alignment.
-Specifically, FedMAE’s masking strategies might hurt the alignment of image and text modalities, and FedX’s single-modality strategy can’t be adapted to model the image-text cross-modal interaction.
-Instead, we think FedAvg is still a strong baseline compared to other adapted methods, as we observed in the experiments.
-That’s because FedAvg doesn’t hurt the ability to learn cross-modal alignment of backbone pre-training models by adding augmented losses.
-Our method can be transferred to multiple multi-modal pre-training methods.
-Table 3 shows the downstream task performance of the MGCA and the GLoRIA backbone pre-training method when combined with our strategy.
-Our method has successfully upgraded MGCA and GLoRIA to adapt to the heterogeneous federated multi-modal pre-training scene, as shown in the significant improvement of the classification and segmentation downstream tasks.
-5.4 Analysis experiment
-In this section, we mainly conduct analysis experiments using the SLIA backbone method for pre-training unless specifically stated.
-Table 4: Ablation results. We report the few-shot classification accuracy and the retrieval accuracy.
-Strategy $1\%$ (cls.) Chest imaG. (ret.)
-Covid RSNA Rec.@1 Rec.@5 Wst.@1 Wst.@5
-[…]
-[…]
-ours 83.21 80.50 29.49 26.99 73.03 67.65
-Table 5: Ablation experiment results on the zero-shot generalization ability of our method on the real-world dataset.
-Strategy O.O.D In domain In domain Wst.
-Rec.@1 Rec.@5 Rec.@1 Rec.@5 Rec.@1 Rec.@5
-FedX 12.31 61.04 3.22 16.76 1.58 11.94
-FedMAE 12.93 63.72 5.05 23.91 4.34 20.86
-FedEMA 15.19 66.94 24.01 67.05 21.85 62.44
-FedAvg 15.3 66.62 28.83 72.12 25.34 66.68
-FedRGB (Ours) 15.43 67.29 29.49 73.03 26.99 67.65
-Figure 8: The trade-off adjusted by the hyper-parameter $\alpha$.
-Figure 9: The role of the hyper-parameter $\alpha$.
-Figure 10: The role of the radius of the uncertainty set $\rho$.
-Robust teacher align module is better than simply averaged one.
-To investigate the efficiency of the robust align module, we replace it with an align module that is simply averaged by local ones and pre-train a new model.
-As shown in table 4, the downstream performance of the replaced one is lower than the original one.
-This means the robust align module can enhance the ability to capture cross-modal alignments and learn transferable representation.
-Strong regularization is better than weak regularization for guiding the updation.
-To investigate the effect of the guidance from the robust align module, we replace the L-2 regularization with contrastive loss as a weaker regularization and pre-train a new model.
-As shown in table 4, the weaker guidance leads to lower classification accuracy and retrieval accuracy in the downstream tasks.
-Guidance hyper-parameter adjusts the average and disparity of client performances.
-Figure 8 shows that higher guidance achieves smaller disparity while bringing less harm on average performance.
-We changed the hyperparameter $\alpha$ from 2 to 10, the disparity of the retrieval accuracy on test sets of different clients decreased significantly while there was only a slight loss on the averaged performance.
-On the other hand, Figure 9 shows that increasing $\alpha$ wouldn’t result in a penalty in the transferability of the pre-trained backbone, the downstream task performances are maintained at a high level as we enlarge the guidance.
-Detailed results can be seen in Table 15.
-Figure 11: Robustness & convergence on the number of com. turns.
-Convergence check: The downstream performance of the pre-trained got promoted as we increased the number of comm.
-turns.
-After 20 communications, the performance turned out to be stable which verified the convergent ability of our method.
-Detailed results can be seen in Table 14.
-Our method has better classification accuracies on unseen out-of-distribution multi-modal X-Ray datasets.
-The multi-modal retrieval accuracy in the unseen domain of our method beats the baseline by a large margin as shown in Table 5, which verifies our method’s ability to learn generalizable cross-modal alignment.
-Meanwhile, our FedRGB has higher in-domain averaged and worst-case acc..
-Therefore, improvements in the in-domain performance of FedRGB do not come at the cost of decreased out-of-domain performance.
-Smaller radius of uncertain set brings higher generalization and transferability on downstream task.
-Figure 10 shows the downstream performance of models pre-trained with different radius of uncertainty set in the DRO process.
-As the $\rho$ increases, the performance of fine-tuning-based downstream tasks shows an upward trend.
-Detailed results can be seen in Table 10.
-6 Conclusion
-Data limitation is a long-standing problem in the multi-modal learning domain.
-Despite federated learning can leveraging datasets from multiple sources while guaranteeing privacy issues, its performance would be damaged by data heterogeneity.
-Inspired by our empirical findings on the impact of heterogeneity on federated multi-modal learning, we propose the FedRGB framework to multigate heterogeneity for federated biomedical vision-language pre-training.
-The efficiency of our method has been verified by comprehensive experiments.
-7 Limitation
-The design of the proposed FedRGB approach has taken it into serious consideration to mitigate the potential negative societal impact..
-For example, to tackle the data privacy issue, FedRGB has avoided the transmission of any data or latent representations during all server-client communications, while only transferring model parameters and loss, thereby minimizing the risk of privacy leakage in federated learning as much as possible.
-However, our model may still be subject to privacy-related issues due to the communication of model parameters, which is a common problem faced by other federated learning methods as well.
-There are also research studies in this line to avoid privacy leakage from model parameters.
-For fairness, our method aims at optimizing the performance of the worst-case client to promote model equity across different cohorts, whichhas shown great generalizable ability on unseen data.
-This approach ensures that the pre-trained model not only performs well and fairly on the data from hospitals participating in the federated training but also demonstrates good performance on hospitals that did not participate in the training.
-Meanwhile we acknowledge that there may still exist limitations.
-Firstly, If we can resolve the trade-off between privacy protection and data representation communication, it may allow us to design more advanced algorithms to further enhance the performance of federated multimodal pre-training models.
-Furthermore, in this work we mainly consider the multimodal medical datasets of X-Ray imageeport due to the limitation of the public datasets.
-However, some rare diseases and more complicated modalities in various medical applications lack paired data, which] may bring other challenges for federated multimodal pretraining.
-8 Acknowledgement
-Zitao Shuai and Liyue Shen acknowledges support from U-M MIDAS PODS Grant and U-M MICDE Catalyst Grant, and computing resource support from NSF ACCESS Program.
-This work used NCSA Delta GPU through allocation CIS230133 and ELE230011 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services &\&& Support (ACCESS) program, which is supported by National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296.
-Appendix A Problem Formulation
-We aim at mitigating the data heterogeneity issues in federally pre-training biomedical vision-language models to better capture the semantic alignment between the two modalities, with heterogeneous private datasets from different clients.
-In this section, we will formulate our problem set-up, assumptions, and related notations.
-Federated Vision-Language Pre-training.
-In this paper, we consider the federated vision-language pre-training (VLP) problem in the biomedical domain.
-Suppose we are given $N$ clients $\{C_{i}\}_{i=1}^{N}$, each client has a corresponding local yet private dataset $\{D_{i}\}_{i=1}^{N}$ that can’t be easily shared due to privacy issues.
-Without access to centralized pre-training, federated pre-training aims to conduct vision-language representation learning across local clients, while still capture a robust and generalizable cross-modal alignment, by efficiently aggregating client models $\{M_{i}\}_{i=1}^{N}$ trained on their private datasets.
-Similar to conventional federated learning scenarios, our federated pre-training consists of multiple communication turns.
-For communication turn $t$, client models $\{M_{i}\}_{i=1}^{N}$ completed the same number of local iteration steps respectively, and the server receives and aggregates the weights sent from local models to produce the aggregated server model $M^{t}$.
-Then it will be sent back to clients and re-trained on $\{D_{i}\}_{i=1}^{N}$ separately to get new local models $\{M_{i}^{t+1}\}_{i=1}^{N}$.
-The federated vision-language pre-training is a more practical and general task compared to related works.
-Table 1 lists the most similar settings in previous works, and the difference between our task and theirs.
-Data Formulation.
-We consider the multimodal datasets including vision and language modalities $X,Y$, consisting of respective data samples $X=\{x_{i}\}_{i=1}^{n},Y=\{y_{i}\}_{i=1}^{m},n,m\in N^{+}$.
-We assume there exists latent variables U that embed semantics across two modalities such that $X=T_{x}(U),Y=T_{y}(U)$, where $T_{x}(\cdot),T_{y}(\cdot)$ are assumed transformations that is not explicit to know.
-For example, the disease category of a given X-ray radiology-report pair is a latent variable connecting these two modalities, which determines the pathology region of the radiology image and semantics in the diagnosis report.
-Thus, we assume each client has paired image-text data that are implicitly correlated through the same latent variables.
-Therefore, local client datasets could be written as $\mathcal{D}:\{D_{i}:(T_{x}(u),T_{y}(u)),u\sim P_{i}(U)\}_{i=0}^{N}$.
-Here, we assume the local datasets are heterogeneous, which means the distribution of latent variables $P_{i}(U)$ is different across clients.
-We formulate two distributions $\mu_{i}\in P(X_{i}),\nu_{i}\in P(Y_{i})$ following (Chen et al., 2020a) as $\mu_{i}=\sum_{j=1}^{n_{i}}a_{i}\delta_{T_{x}(u_{j}))_{i,j}}$ and $\nu_{i}=\sum_{j=1}^{n_{i}}b_{j}\delta_{T_{y}(u_{j})_{i,j}}$, with $\delta_{x}$ as the Dirac function centered on $x$.
-The weight vectors $a_{i}=\{a_{j}\}_{j=1}^{n_{i}}\in\Delta_{n_{i}}$ and $b_{i}=\{b_{j}\}_{j=1}^{n_{i}}\in\Delta_{n_{i}}$ belong to the $(n-1)$- dim simplex, where $\mu_{i}$ and $\nu_{i}$ are probability distributions.
-This formulation can also be applied to feature vectors or representations.
-For a given set of distributions $\{P_{i}\}_{i=1}^{N}$, we can form a new distribution $P^{c}=\sum_{i=1}^{N}w_{i}P_{i},\textbf{w}\in\Delta$ where $\Delta$ is a $(N-1)$-dim simplex by weighted averaging the sample vectors.
-Model Architectures.
-Following the VLP works (Radford et al., 2021; Yang et al., 2022), we frame the VLP using a dual-encoder structure, where the model for each modality consists of a feature encoder module $f(\cdot)$ that outputs feature vectors given the raw input, and a feature alignment module $g(\cdot)$ that projects feature vectors to aligned representations in the shared embedding space.
-Denote the feature vector as $s$, and the aligned representation as $z$.
-We model the outputs of the encoder and alignment modules as conditional distributions.
-Specifically, the encoder takes the multi-modal raw input to generate semantic embeddings of the corresponding modality, which could be formulated as $P(s|U)$.
-Similarly, that of the alignment module can be formulated as $P(z|S,U)$.
-Appendix B Details of Learning Unbiased Cross-modal Alignment via Distribution-based Optimization
-In order to guide the local client training, we need to provide a teacher alignment module as mentioned in the last section, which can capture globally unbiased and generalizable cross-domain alignment.
-How can we obtain such a teacher guidance model?
-We propose to learn the unbiased cross-domain alignment by considering the robustness across various distributions of local clients.
-A teacher alignment module should avoid distorting the aggregated feature encoder while learning the cross-modal alignments.
-Additionally, we introduce a much harder task to force the module to be more robust to biased alignment, by considering the influence of the biased alignment module on the feature encoder.
-Furthermore, to learn a teacher model that performs well on varies data distributions, we minimize distortion degree under the most biased local alignment with the “worst-case” local data distribution.
-Inspired by the distributionally robust optimization (DRO) (Cotter et al., 2019), we model the de-centralized pre-training framework as a two-stage DRO learning process.
-In the first stage, the local models are trained with $\mathcal{L}_{local}$ (Eq.
-2), while the whole model is jointly optimized with the guidance of the teacher alignment module.
-In the second stage, we learn a robust teacher alignment module that performs well conditioned on the worst-case local models where the feature encoder is distorted the most severely.
-We aim to optimize the alignment module on the worst-case distribution of an uncertainty set constructed by the local aligner and local data distribution, which represents possible cases that the model would encounter.
-We construct the uncertainty set based on the distribution of client datasets and local aligners learned in the first stage.
-From the probabilistic view, the local aligners could be written as $P(z|U,f)$, which takes the output of the given encoder $f$ to generate the aligned representations, while $f$ takes the raw input from the local data distribution $P(U)$ to generate the semantic embedding that has not well learned the cross-modal alignment.
-Therefore, the distribution to be considered should be $P(z,U|f)$.
-Hence the uncertainty set can be written as:
-$\footnotesize Q^{c}:\{D(Q||P^{c})\leq\rho\},P^{c}=\sum_{i=1}^{N}w_{i}P_{i}(z,U|f),w\in\Delta_{N-1}$ (7)
-where $N$ is the number of clients, $\Delta$ is a $(N-1)$-dim simplex, $D$ is a discrepancy measure, and $\rho$ is the uncertainty radius that measures the uncertainty degree of the unseen distributions and influences model’s generalization ability.
-Finally, with the teacher align module denoted as $g^{*}$, indicating the aggregated encoder after the first stage training, the learning object in the second stage can be formulated as:
-$\footnotesize min_{g^{*}}\ max_{P\in Q^{c}}\ \mathcal{L}_{dro}(g^{*},P,f^{*})$ (8)
-Following (Zhang et al., 2023b), the optimization goal can be constructed by distributions $P_{i}$ of the support set:
-$\footnotesize min_{g^{*}}\ max_{\textbf{w}\in\Delta_{N-1}}\sum_{i=1}^{N}w_{i}\ \mathcal{L}_{dro}(g^{*},P_{i},f),s.t.D(N\textbf{w}||\textbf{1})\leq\rho$ (9)
-where $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$ are the losses calculated on the local clients respectively.
-While we learn $g^{*}$ by gradient descent, the mirror gradient descent process of optimizing w could be solved efficiently (Zhang et al., 2023b; Duchi et al., 2008):
-$\footnotesize{w_{i}}=\text{Proj}_{\mathcal{Q}{c}}\left({w_{i}e^{\gamma v_{i}}}/\sum_{i=1}^{N}{w_{i}e^{\gamma v_{i}}},\rho\right)$ (10)
-where $\text{Proj}_{\mathcal{Q}{c}}$ refers to the projection to $\{w:D(N\textbf{w}||\textbf{1})\leq\rho\}$, $v_{i}$ represents the loss $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$.
-We alternatively optimize the w and the teacher align module.
-For each communication turn of the federated learning, we first obtain an averaged encoder, and a set of local align modules in the first stage.
-Then we solve the weights and optimize the parameters of the teacher align module.
-Hence the w could be written in a more trackable form as:
-$\footnotesize{w_{i}^{(r+1)}}=\text{Proj}{\mathcal{Q}{c}}\left({w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}}/\sum_{i=1}^{N}{w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}},\rho\right)$ (11)
-where $r$ refers to the communication turn..
-Finally, we formulate $\mathcal{L}_{dro}$ to measure the performance of the teacher align module of alleviating distortion on the encoder.
-Based on equations (3) and (4), we update the teacher alignment module in local clients with weights w.
-Specifically, given a client $c$, we use $f^{*}$ learned in the first stage of the current communication turn to generate features $(\tilde{v^{c}},\tilde{l^{c}})$, and a learnable copy of $f^{*}$ to generate distorted features $(v^{c},l^{c})$.
-The distorted features are input to the frozen $g$ to calculate the task loss $L_{task}$, and are input to the $g^{*}$ to get calibrated representations $(z_{I}^{c*},z_{T}^{c*})$.
-Meanwhile, $(\tilde{v^{c}},\tilde{l^{c}})$ are input to $g^{*}$ to get representation $\tilde{(z_{I}^{c*}},\tilde{z_{T}^{c*}})$ for regularization.
-The regularization term could be written as:
-$\footnotesize\mathcal{L}_{dro,\tau}=\beta(\sum_{i=1}^{N_{c}}(\tilde{z^{c*}}_{I,i}-z^{c*}_{I,i})^{2}+\sum_{i=1}^{N_{c}}(\tilde{z^{c*}}_{T,i}-z^{c*}_{T,i})^{2})+\mathcal{L}_{CL,\tau}$ (12)
-where $\mathcal{L}_{CL}$ is the contrastive loss (Radford et al., 2021) between $(z^{fz,c*}_{img,i},z^{fz,c*}_{txt,i})_{i=0}^{N_{c}}$, and $N_{c}$ refers to the number of samples in the client $c$.
-Appendix C Experiment Details
-C.1 Pre-training setup
-Following (Wang et al., 2022) we utilize the MIMIC-CXR (Bigolin Lanfredi et al., 2022) dataset for multi-modal pre-training.
-This dataset is widely used in the biomedical multi-modal learning domain, with $227,835$ image-text pairs from $65,379$ patients.
+Thus, we are motivated to propose our FedAID.
+Appendix B Experiment Details
+B.1 Pre-training setup
+Following Wang et al.
+(2022) we utilize the MIMIC-CXR Bigolin Lanfredi et al.
+(2022) dataset for multi-modal pre-training.
+This dataset is widely used in the medical multi-modal learning domain, with $227,835$ image-text pairs from $65,379$ patients.
 Some related works also have imported additional features to image-text pairs to augment the data.
 However, we only use the image-text pairs for pre-training to make the results more generalizable.
+The MIMIC-CXR dataset is open access, it can be obtained through MIMIC-CXR Access.
 During the pre-training, local clients only have access to their highly heterogeneous datasets.
-To construct the heterogeneous client datasets, following (Yan et al., 2023) we employ the Latent Dirichlet Allocation (LDA) (Blei et al., 2003) to divide the MIMIC-CXR dataset into $5$ partitions based on a selected sensitive attribute.
-For implementation, we import the corresponding attribute information of given image-text pairs from the Chest-Imagenone dataset (Wu et al., 2021) and divide local datasets based on disease category.
+To construct the heterogeneous client datasets, following Yan et al.
+(2023) we employ the Latent Dirichlet Allocation (LDA) Blei et al.
+(2003) to divide the MIMIC-CXR dataset into $5$ partitions based on a selected sensitive attribute.
+For implementation, we import the corresponding attribute information of given image-text pairs from the Chest-Imagenone dataset Wu et al.
+(2021) and divide local datasets based on disease category.
 The disease category is a multi-label binary attribute and is transformed into a multi-class label.
-That’s because the words in the clinical report are highly related to the disease category as illustrated in 12.
+That’s because the words in the clinical report are highly related to the disease category as illustrated in Fig 9.
 We set the heterogeneity degree in the LDA algorithm to be 1 and the subgroups of the MIMIC-CXR that are used as local datasets are summarized in Table 4.
-Specifically, we select 5 commonly considered diseases (Bannur et al., 2023): ’Edema’,’Pleural Effusion’, ’Consolidation’, ’Pneumothorax’, and ’Pneumonia’.
+Specifically, we select 5 commonly considered diseases Bannur et al.
+(2023): ’Edema’,’Pleural Effusion’, ’Consolidation’, ’Pneumothorax’, and ’Pneumonia’.
 We set the non-NaN value to 1 and then set NaN value to 0 to construct a 5-way binary multi-label.
@@
 The distribution of the disease of the client dataset is summarized in Table 4.
-Figure 12: Illustration of the divided heterogeneous subgroups for each client.
-Table 6: Distortion of local training of FedAvg method.
-We show the performance gap between the server and local retrained models.
-We employ the decrease in the similarity of the image embedding and text embedding to be an indicator of distortion.
-FedAvg Test on C1 Test on C2 Test on C3 Test on C4 Test on C5
-Retrain On C1 0.092 0.078 0.08 0.073 0.082
-Retrain On C2 0.054 0.049 0.046 0.047 0.048
-Retrain On C3 0.069 0.072 0.07 0.066 0.067
-Retrain On C4 0.092 0.09 0.084 0.093 0.088
-Retrain On C5 0.073 0.08 0.07 0.068 0.087
-Avg 0.076 0.0738 0.07 0.0694 0.0744
+Figure 9: Illustration of the strong connection between latent variable and the text modality.
 Table 4: The distribution of the sensitive attribute age-decile. From this table, we can verify that the local datasets are highly heterogeneous.
@@
 We divide the MIMIC-CXR into 5 heterogeneous subgroups to construct 5 client datasets.
-Each divided dataset consists of train splits and validation splits based on the notation of the MIMIC-CXR.
+Each divided dataset consists of train splits and test splits based on the notation of the MIMIC-CXR.
 We use $404\times A$ 40 for pre-training and set the batch size to 388.
 We set the learning rate to $2\times 10^{-5}$, the number of communications to 25.
-We set the updating step size $\gamma=1$, and uncertainty radius $\rho=1$, hyper-parameter $\alpha=1,\beta=2$ in the main experiments.
+We set the uncertainty radius $\rho=1$, hyper-parameter $\alpha=1,\beta=2$ in the main experiments.
 For each communication, we randomly sample 50 batches of data from the client datasets.
-To obtain the guidance, we run an extra communication for our method.
-C.2 Downstream tasks
+B.2 Downstream tasks
 We conduct Few-shot classification, Medical image segmentation, and Image Retrieval downstream tasks to evaluate the generalization ability of the pre-trained model.
 Few-shot classification.
-To verify the efficiency of the pre-trained model on the general medical image task, we test their performance on multiple image classification benchmarks: (1) RSNA Pneumonia Detection (RSNA) (Shih et al., 2019), the task is to predict if the image reflects a pneumonia-positive phenomenon.
-(2) Covidx (Wang et al., 2020), which has three categories that COVID-19, non-COVID pneumonia and normal.
+To verify the efficiency of the pre-trained model on the general medical image task, we test their performance on multiple image classification benchmarks: (1) RSNA Pneumonia Detection (RSNA) Shih et al.
+(2019), the task is to predict if the image reflects a pneumonia-positive phenomenon.
+(2) Covidx Wang et al.
+(2020), which has three categories that COVID-19, non-COVID pneumonia and normal.
 We fine-tune our pre-trained model with an additional linear layer on 1%, 10% percentage of the training dataset of downstream tasks, and evaluate the image classification accuracy.
@@
 Medical image segmentation.
-To investigate the transfer ability of our model on fine-grained tasks, we conduct experiments on medical image segmentation on the RSNA (Wang et al., 2020) benchmark.
-We follow (Wang et al., 2022) to convert object detection ground truths of RSNA into masks for semantic segmentation.
-Similar to (Huang et al., 2021), we construct a U-Net framework using a decoder and our pre-trained image encoder.
+To investigate the transfer ability of our model on fine-grained tasks, we conduct experiments on medical image segmentation on the RSNA Wang et al.
+(2020) benchmark.
+We follow Wang et al.
+(2022) to convert object detection ground truths of RSNA into masks for semantic segmentation.
+Similar to Huang et al.
+(2021a), we construct a U-Net framework using a decoder and our pre-trained image encoder.
 During the fine-tuning, we freeze the encoder and fine-tune the decoder using 1%, 10% of the training data of the downstream dataset, and then use the Dice score to evaluate the performance of our model.
@@
 Out-of-Distribution testing benchmark.
-We adapt the EHRXQA (Bae et al., 2023) VQA dataset for evaluation.
-This dataset is built on the meta-data notation of MIMIC-CXR (Bigolin Lanfredi et al., 2022) and ChestImaGenome (Wu et al., 2021).
+We adapt the EHRXQA Bae et al.
+(2023) VQA dataset for evaluation.
+This dataset is built on the meta-data notation of MIMIC-CXR Bigolin Lanfredi et al.
+(2022) and ChestImaGenome Wu et al.
+(2021).
 We merge the question and answer sentences as the text input and use the corresponding image to construct the image-text pairs for the multi-modal retrieval task.
 Other settings are the same as the ordinary image retrieval task.
-C.3 Backbones and Baselines
-We focus on upgrading biomedical vision-language pre-training (VLP) strategies to heterogeneous federated learning scenes.
-Since federated heterogeneous VLP is a new topic, we adopt existing biomedical VLP baselines to federated strategies (Lu et al., 2023a).
-Specifically, we consider adopting simple language-image contrastive alignment (ConVIRT) (Zhang et al., 2022b; Radford et al., 2021), global-local language-image contrastive alignment(GLoRIA) (Huang et al., 2021), and Multi-Granularity Cross-modal Alignment (MGCA) (Wang et al., 2022).
-For federated learning strategies, we consider simple averaging (FedAvg) (McMahan et al., 2017; Lu et al., 2023a), centralized training (Local), and decentralized training (Global).
-For fair comparisons, we combine each federated learning strategy with different biomedical VLP methods.
+B.3 Backbones and Baselines
+We focus on enabling medical vision-language pre-training (VLP) methods to be applied to heterogeneous federated learning scenes and have considered the generalization ability of FedAID on different backbone VLP methods.
+Since federated heterogeneous VLP is a new topic, we adopt existing medical VLP baselines to federated strategies Lu et al.
+(2023a).
+Specifically, we consider adopting contrastive learning based methods: simple language-image contrastive alignment (ConVIRT) Zhang et al.
+(2022b); Radford et al.
+(2021), global-local language-image contrastive alignment(GLoRIA) Huang et al.
+(2021a), and Multi-Granularity Cross-modal Alignment (MGCA) Wang et al.
+(2022).
+Since the federated multi-modal per-training is a novel setting, we adapted FedMAE Yan et al.
+(2023), FedEMA Zhuang et al.
+(2022), FedMOON Li et al.
+(2021b) for comparison.
+These are self-supervised learning methods in the federated learning domain which also focus on tackling the data heterogeneity.
+For basic federated learning baselines, we consider simple averaging (FedAvg) McMahan et al.
+(2017); Lu et al.
+(2023a), decentralized training (Local), and centralized training (Global).
+For fair comparisons, we combine each federated learning strategy with different medical VLP methods.
 For baselines pre-trained in Local strategy, we report the averaged performance of the local models.
@@
 We utilize a ViT-Base model as our vision encoder and utilize the Bert-Base as our text encoder.
-We pre-process the input image following (Wang et al., 2022) to resize the radiology graph to the size of $224\times 224$ to match the vision encoder.
+We pre-process the input image following Wang et al.
+(2022) to resize the radiology graph to the size of $224\times 224$ to match the vision encoder.
 Table 5: Descriptive Statistics of the subgroups.
-Disease C1 C2 C3 C4 C5
+Disease Client 1 Client 2 Client 3 Client 4 Client 5
 Edema 13150 8061 13336 8121 8341
-P.E. 17048 20865 17706 9974 10424
-Consolid. 4879 2201 2221 4024 4615
-Pneumoth. 14177 6962 6315 13991 7115
+Pleural Effusion. 17048 20865 17706 9974 10424
+Consolidation 4879 2201 2221 4024 4615
+Pneumothorax 14177 6962 6315 13991 7115
 Pneumonia 11605 10105 7152 3431 7064
-Appendix D Additional Analysis Experiment
-We report the communication costs and training costs of our method and other FL methods on both the server and client sides in Table 9.
-As shown in the table, the additional computational overhead introduced by our method is within an acceptable range.
-Table 9: Summarization of computation costs.
-Method Train Params Train time Server Cost Info. Trans.
-De-centralized 172.6 M 14.62s/it 863M FLOP 172.6 M
-FedAvg 172.6 M 14.62s/it 863M FLOP 172.6 M
-FedMAE 226.1M 14.95s/it 1130.5M FLOP 226.1M
-FedX 172.6 M 22.66s/it 863M FLOP 172.6 M
-FedEMA 172.6 M 14.59s/it 863M FLOP 172.6 M
-Ours [1st stage] 190.4M 15.4s/it 951.2M FLOP 190.4M
-Ours [2nd stage] 190.4M 23.68s/it 89M FLOP 17.8M
-Centralized 172.6 M 14.62s/it 863M FLOP 172.6 M
+Figure 10: Role of guidance degree $\alpha$. Increasing $\alpha$ results in at most small penalty in the transferability.
+Figure 11: Role of the uncertainty radius $\rho$. A small radius of $\rho$ brings higher generalization and transferability on downstream tasks.
+Figure 12: Robustness & convergence on the number of com. turns.
+Appendix C Additional Analysis Experiment
+Table 6: Distortion of local training of FedAvg method.
+We show the performance gap between the server and local retrained models.
+We employ the decrease in the similarity of the image embedding and text embedding to be an indicator of distortion.
+FedAvg Test on C1 Test on C2 Test on C3 Test on C4 Test on C5
+Retrain On C1 0.092 0.078 0.08 0.073 0.082
+Retrain On C2 0.054 0.049 0.046 0.047 0.048
+Retrain On C3 0.069 0.072 0.07 0.066 0.067
+Retrain On C4 0.092 0.09 0.084 0.093 0.088
+Retrain On C5 0.073 0.08 0.07 0.068 0.087
+Avg 0.076 0.0738 0.07 0.0694 0.0744
 Table 7: Detailed results on the analysis experiments of the role of hyper-parameter $\alpha$.
@@
 $\rho=1$ 83.16 83.41 80.00 89.75 66.98 70.18
-Our method has succesfully avoided the distortion caused by heterogeneous datasets during the local training.
-Empirically, we demonstrate the effectiveness of this guidance-based local training strategy.
-Specifically, we retrain the model on each local dataset with the same backbone methods and setting and conduct the study on the change of similarity between embeddings of paired image and text.
-We show the performance gap between the server and local retrained models.
-The reduced similarity of the image embedding and text embedding from paired data can be regarded as an indicator of ”distortion” of cross-modal alignment.
-As reported in Table 6 and Table 11, we observe that the average decrease on the similarity of image and text embedding of our method is much smaller compared to FedAvg baseline.
-Role of the number of clients that participated in the federated pre-training.
-We have conducted experiments on how the number of clients participating in the federated pre-training affects the performance of the learned model.
-As shown in the Table 16, we have observed that bringing more diverse clients for the federated pre-training is likely to improve the performance of the pre-trained model.
-Appendix E Detailed Experiment Results
-Here we show the detailed experiment results.
+Table 8: Ablation experiment results on the zero-shot generalization ability of our method on the real-world dataset.
+Strategy O.O.D In domain In domain Wst.
+Rec.@1 Rec.@5 Rec.@1 Rec.@5 Rec.@1 Rec.@5
+FedMAE 12.93 63.72 5.05 23.91 4.34 20.86
+FedEMA 15.19 66.94 24.01 67.05 21.85 62.44
+FedAvg 15.3 66.62 28.83 72.12 25.34 66.68
+FedAID (Ours) 15.43 67.29 29.49 73.03 26.99 67.65
 Table 9: Distortion of local training of ours method.
@@
 Avg 0.0508 0.0452 0.038 0.0506 0.039
-[…]
+Our method has succesfully avoided the distortion caused by heterogeneous datasets during the local training.
+Empirically, we demonstrate the effectiveness of this guidance-based local training strategy.
+Specifically, we retrain the model on each local dataset with the same backbone methods and setting and conduct the study on the change of similarity between embeddings of paired image and text.
+We show the performance gap between the server and local retrained models.
+The reduced similarity of the image embedding and text embedding from paired data can be regarded as an indicator of "distortion" of cross-modal alignment.
+As reported in Table 6 and Table 9, we observe that the average decrease on the similarity of image and text embedding of our method is much smaller compared to FedAvg baseline.
+Role of the number of clients that participated in the federated pre-training.
+We have conducted experiments on how the number of clients participating in the federated pre-training affects the performance of the learned model.
+As shown in the Table 14, we have observed that bringing more diverse clients for the federated pre-training is likely to improve the performance of the pre-trained model.
+Our method has better classification accuracies on unseen out-of-distribution multi-modal X-Ray datasets.
+The multi-modal retrieval accuracy in the unseen domain of our method beats the baseline by a large margin as shown in Table 8, which verifies our method’s ability to learn generalizable cross-modal alignment.
+Meanwhile, our FedAID has higher in-domain averaged and worst-case acc..
+Therefore, improvements in the in-domain performance of FedAID do not come at the cost of decreased out-of-domain performance.
+Convergence check: As shown in Figure 12 the downstream performance of the pre-trained got promoted as we increased the number of comm.
+turns.
+After 20 communications, the performance turned out to be stable which verified the convergent ability of our method.
+Detailed results can be seen in Table 12.
+Appendix D Detailed Experiment Results
+In addition to the experiment results shown in the above figures and tables, here we show the detailed points in Table 9, Table 10, Table 11, Table 15, Table 16, Table 17, Table 12, Table 13 and Table 14.
+Table 10: The comparison of retrieval acc.
+on each client denoted as $\{C_{i}\}_{i=1}^{5}$,of centralized, FedAvg, and averaged acc.
+of decentralized pre-trained models using the ConVIRT backbone.
+We report the local models of decentralized pre-training strategy as Decentralizedi.
 Strategy Recall@1 (ACC) Recall@5 (ACC)
@@
 FedAvg 30.43 25.34 26.90 28.75 32.73 28.83 76.39 66.68 69.75 73.84 73.94 72.12
-[…]
-[…]
-[…]
-[…]
-[…]
+Decentralized1 17.68 14.71 15.37 18.18 14.14 16.01 56.99 49.56 51.29 54.55 55.10 53.50
+Decentralized2 15.33 11.62 13.85 14.96 13.83 13.92 54.84 41.93 46.20 47.90 45.62 47.30
+Decentralized3 17.41 13.08 14.12 15.07 15.39 15.01 50.40 44.21 46.37 49.72 52.53 48.65
+Decentralized4 16.54 14.34 14.08 15.36 17.35 15.53 57.02 45.72 47.45 52.38 57.63 52.04
+Decentralized5 21.68 14.34 14.20 15.56 19.50 17.06 57.89 48.43 50.18 52.36 61.66 54.10
 Local.avg. 17.73 13.62 14.32 15.83 16.04 15.51 55.43 45.97 48.30 51.38 51.12 51.12
-[…]
+Table 11: The performance of the server model after 25 commu.
+turns and the averaged performance of corresponding local models after 25 and 26 commu.
+turns, on each client denoted as $\{C_{i}\}_{i=1}^{5}$.
+We utilize the ConVIRT as the backbone.
+We report the models after local update on client datasets of FedAvg pre-training strategy as Locali.
 Strategy com. turn Recall@1 (ACC) Recall@5 (ACC)
@@
 FedAvg 25 30.43 25.34 26.90 28.75 32.73 76.39 66.68 69.75 73.84 73.94
-[…]
-[…]
-[…]
-[…]
-[…]
+Local0 25 31.75 23.37 25.96 26.21 33.71 73.44 64.14 67.15 71.38 71.79
+Local1 25 27.72 22.52 23.77 25.06 27.70 73.45 63.08 64.57 69.15 68.58
+Local2 25 28.55 24.37 24.30 27.26 28.93 73.26 64.94 67.52 70.28 71.75
+Local3 25 30.63 22.62 23.90 25.36 26.42 72.57 64.02 66.26 68.92 67.62
+Local4 25 27.31 24.37 25.47 26.47 28.93 73.26 65.83 68.95 71.23 69.21
 1-5 Avg. 25 29.19$\downarrow$ 23.45$\downarrow$ 24.68$\downarrow$ 26.07$\downarrow$ 29.14$\downarrow$ 73.20$\downarrow$ 64.40$\downarrow$ 66.89$\downarrow$ 70.20$\downarrow$ 69.79$\downarrow$
-[…]
-[…]
-[…]
-[…]
-[…]
+Local0 26 29.91 23.18 25.24 26.66 33.05 73.44 64.14 67.15 71.38 71.79
+Local1 26 28.7 23.05 23.95 25.91 27.37 74.48 63.37 64.12 69.69 67.93
+Local2 26 30.9 23.85 25.42 27.67 29.89 72.40 65.57 67.42 71.56 71.12
+Local3 26 30.11 22.66 23.41 24.45 27.05 73.09 63.48 65.32 67.48 67.31
+Local4 26 26.96 24.81 25.47 26.6 29.55 73.61 66.03 69.17 71.19 69.54
 1-5 Avg. 26 29.32$\downarrow$ 23.51$\downarrow$ 24.70$\downarrow$ 26.26$\downarrow$ 29.38$\downarrow$ 73.33$\downarrow$ 64.44$\downarrow$ 66.34$\downarrow$ 70.10$\downarrow$ 69.41$\downarrow$
@@
 n=5 83.21 83.54 80.5 90 71.51 74.08 29.49 73.03 26.99 67.65
-[…]
+Table 15: The accuracy of the test set of each client.
+We show the performance of FedAvg pre-trained baseline and its retrained models on different client datasets.
+We report the models after local update on client datasets of FedAvg pre-training strategy as Locali.
 position model com. Recall@1 (ACC) Recall@5 (ACC)
@@
 - server 50 32.29 26.03 26.99 27.06 30.23 28.52$\downarrow$ 77.60 67.86 69.43 72.06 71.74 71.74$\downarrow$
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+$\rightarrow$ shallow Local0 25 30.35 24.95 25.32 28.44 28.62 27.54$\downarrow$ 73.78 67.22 68.44 71.95 73.01 70.88$\downarrow$
+$\rightarrow$ shallow Local1 25 34.25 26.4 27.27 29.66 30.20 29.44$\uparrow$ 78.30 69.75 72.33 75.07 77.42 74.57$\uparrow$
+$\rightarrow$ shallow Local2 25 33.68 26.39 27.27 29.66 30.20 29.44$\uparrow$ 77.25 67.37 70.67 74.21 70.51 72.02$\uparrow$
+$\rightarrow$ shallow Local3 25 27.65 18.94 19.30 25.61 24.87 24.95$\downarrow$ 72.40 64.19 64.82 70.47 71.13 68.60$\downarrow$
+$\rightarrow$ shallow Local4 25 26.19 18.94 19.30 22.74 22.66 21.97$\downarrow$ 69.27 56.76 58.94 63.35 64.50 62.56$\downarrow$
+Table 16: The accuracy of the model on each client.
+We show the acc.
+of centralized and FedAvg pre-trained baselines and de-centralized pre-trained models shown as Locali retrained on the union of training splits of client datasets.
+We fine-tune shallow layers of the de-centralized pre-trained model with the union dataset.
 strategy model Recall@1 (ACC) Recall@5 (ACC)
@@
 FedAvg server 30.43 25.34 26.90 28.75 32.73 28.83 76.39 66.68 69.75 73.84 73.94 72.12
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Decentralized Local1 28.72 22.61 23.52 22.32 24.59 24.35 70.94 60.77 63.31 62.40 63.42 64.17
+Decentralized Local2 17.36 19.85 18.23 17.58 17.68 18.14 52.70 56.05 55.14 54.01 53.54 54.29
+Decentralized Local3 20.92 20.67 25.96 21.04 22.34 22.19 58.93 58.05 65.34 58.21 58.99 59.90
+Decentralized Local4 20.86 20.07 20.69 25.57 21.12 21.66 59.11 56.88 57.82 64.70 58.69 59.44
+Decentralized Local5 21.79 19.54 21.98 20.90 31.54 23.15 60.73 56.99 59.80 60.17 74.07 62.35
+Decentralized Local1 17.68 14.71 15.37 18.18 14.14 16.02 56.99 49.56 51.29 54.55 55.10 53.50
+Decentralized Local2 15.33 11.62 13.85 14.96 13.83 13.92 54.84 41.93 46.20 47.90 45.62 47.30
+Decentralized Local3 17.41 13.08 14.12 15.07 15.39 15.01 50.40 44.21 46.37 49.72 52.53 48.65
+Decentralized Local4 16.54 14.34 14.08 15.36 17.35 15.53 57.02 45.72 47.45 52.38 57.63 52.04
+Decentralized Local5 21.68 14.34 14.20 15.56 19.50 17.06 57.89 48.43 50.18 52.36 61.66 54.10
 Table 17: Detailed ablation results.
@@
 ours 32.53 26.99 27.70 29.09 31.12 29.62 77.08 67.65 70.32 73.96 76.12 73.14
+Table 18: Comparison between related tasks and our federated multi-modal pre-training task.
+We organize the related works based on our key areas of interest.
+(1).
+Heterogeneous clients.
+(2).
+Client-side pretraining.
+(3).
+Federated self-supervised learning (FSSL) to leverage unlabeled data.
+(4).
+multi-modal alignment.
+(5).
+Pre-train models that can be transferred to other downstream tasks.
+(6).
+No communication on public data or feature representations.
+Task hetero. pretrain on client FSSL. align modalities transfer. no feat. commu.
+Learn different client modalities Chen & Zhang (2022) $\checkmark$ $\checkmark$
+Augment server with local modalityYu et al. (2023) $\checkmark$ $\checkmark$ $\checkmark$
+FSSL. on unlabeled datasets Zhang et al. (2023a) $\checkmark$ $\checkmark$
+FSSL with server fine-tune Yan et al. (2023) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Federated multi-modal pre-training $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Appendix E Problem Position
+The federated vision-language pre-training is a more practical and general task compared to related works.
+Table 18 lists the most similar settings in previous works, and the difference between our task and theirs.
+Our federated multi-modal pre-training focuses on learning generalizable vision-language models which can be transferred to various downstream tasks, with paired unlabeled multi-modal data, in a fully decentralized way that can guarantee the data privacy required by clinical applications.
+We tackle the data heterogeneity problem in local training, which is a vital concern in federated learning.
+During client-server communication, we avoid data transmission on features of local data to ensure data privacy.
+Appendix F Detailed Problem Formulation
+We aim at mitigating the data heterogeneity issues in federally pre-training medical vision-language models from heterogeneous private client datasets.
+In this section, we will formulate the set-up of the Federated Vision-Language Pre-training task, which aims to capture better cross-modal alignment by leveraging diverse semantics from local datasets.
+Then we will introduce our assumption on multi-modal pre-training data and heterogeneity.
+Please see Appendix F for the detailed preliminary.
+Federated Vision-Language Pre-training (FedVLP).
+We aim to federally pre-train medical vision-language models.
+Specifically, given $N$ clients $\{C_{i}\}_{i=1}^{N}$, FedVLP decentralizedly conducts vision-language representation learning across clients to learn local models, and aggregates them on a central server to obtain a more generalizable pre-trained model.
+Each client has a local yet private dataset $\{D_{i}\}_{i=1}^{N}$ that can’t be easily shared due to privacy issues.
+During FedVLP, different from supervised federated learning, we focus on the transferability of the aggregated model.
+Employing an alternative updating approach, federated pre-training also consists of multiple communication turns between clients and the server the same as conventional federated learning (FL) scenes.
+For each turn $t$, client models $\{M_{i}^{t}\}_{i=1}^{N}$ completed local pre-training respectively and send model weights to server, then it receives and aggregates the weights to obtain server model $M^{t}$.
+$M^{t}$ will be broadcast to clients for updating local models, and complete the main part of the communication turn.
+In some FL methods, additional local computation and information transmissions are conducted after clients have received the aggregated model, and they are counted as a part of the communication.
+The federated vision-language pre-training is a more practical and general task compared to related works.
+Table 18 in appendix lists the most similar settings in previous works, and their differences with ours.
+Formulation of Multi-Model Pre-training Data and Heterogeneity.
+We consider the multimodal datasets including vision and language modalities $X,Y$, consisting of respective data samples $X=\{x_{i}\}_{i=1}^{n_{1}},Y=\{y_{i}\}_{i=1}^{n_{2}},n_{1},n_{2}\in N^{+}$.
+We assume there exists discrete mutlti-label latent variables U that embed semantics across two modalities such that $X=H_{I}(U),Y=H_{T}(U)$, where $H_{I}(\cdot),H_{T}(\cdot)$ are assumed transformations that is not explicit to know.
+For example, the disease category of a given X-ray radiology-report pair is a latent variable connecting these two modalities, which determines the pathology region of the radiology image and semantics in the diagnosis report.
+Thus, we assume each client has paired image-text data that are implicitly correlated through the same latent variables.
+Therefore, local client datasets could be written as $\mathcal{D}:\{D_{i}:(H_{I}(u),H_{T}(u)),u\sim P_{i}(U)\}_{i=0}^{N}$, where u can be thought as a latent class.
+Here, we assume the local datasets are heterogeneous, which means the distribution of latent variables $P_{i}(U)$ is different across clients.
+We formulate two distributions $\mu_{i}\in P(X_{i}),\nu_{i}\in P(Y_{i})$ following Chen et al.
+(2020) as $\mu_{i}=\sum_{j=1}^{n_{i}}a_{i}\delta_{T_{x}(u_{j}))_{i,j}}$ and $\nu_{i}=\sum_{j=1}^{n_{i}}b_{j}\delta_{T_{y}(u_{j})_{i,j}}$, with $\delta_{x}$ as the Dirac function centered on $x$.
+The weight vectors $a_{i}=\{a_{j}\}_{j=1}^{n_{i}}\in\Delta_{n_{i}}$ and $b_{i}=\{b_{j}\}_{j=1}^{n_{i}}\in\Delta_{n_{i}}$ belong to the $(n-1)$- dim simplex, where $\mu_{i}$ and $\nu_{i}$ are probability distributions.
+This formulation can also be applied to feature vectors or representations.
+For a given set of distributions $\{P_{i}\}_{i=1}^{N}$, we can form a new distribution $P^{c}=\sum_{i=1}^{N}w_{i}P_{i},\textbf{w}\in\Delta$ where $\Delta$ is a $(N-1)$-dim simplex by weighted averaging the sample vectors.
+Model Architectures.
+Following the VLP works Radford et al.
+(2021); Yang et al.
+(2022), we formulate the VLP using a dual-model structure, where the model for each modality consists of a feature encoder module $f(\cdot)$ that outputs feature vectors given the raw input, and a feature alignment module $g(\cdot)$ that projects feature vectors to aligned representations in the shared embedding space.
+We denote the feature vectors on the text side as $l$, the feature vectors on the image side as $v$, and the aligned representation as $z_{l},z_{v}$ respectively.
+During the distributionally robust optimization process, while we utilize the distribution of latent semantics to construct an uncertainty set for optimization, we take local alignment modules and the server-aggregated encoder as conditions.
+Uncertainty set for Distributionally Robust Optimization (DRO).
+After the first stage in local training, the server aggregates the weights of local encoders to get $f^{*}$ and sends it back to clients for learning a robust teacher alignment module $g^{*}$.
+We aim to optimize the alignment module on the worst-case distribution of an uncertainty set constructed by the local data distribution, which represents possible cases that the model would encounter.
+Specifically, a worst-case distribution is considered based on support distributions and the uncertain radius.
+We first consider the set of possible weighted combinations of support distributions and then enlarge the set by introducing an uncertain radius.
+We utilize distributions of latent semantics $P_{i}(U)$ of local clients as support distributions.
+Therefore, following the notation in F, the uncertainty set can be written as:
+$\footnotesize Q^{c}:\{D(Q||P^{c})\leq\rho\},P^{c}=\sum_{i=1}^{N}w_{i}P_{i}(U),w\in\Delta_{N-1}$ (6)
+where $N$ is the number of clients, $\Delta$ is a $(N-1)$-dim simplex, $D$ is a discrepancy measure, and $\rho$ is the uncertainty radius that measures the uncertainty degree of the unseen distributions and influences model’s generalization ability.
+Appendix G Analysis of Theoretical Insights
+G.1 Theoretical Analysis on Remark 1
+Here we provide a theoretical analysis on Remark 1 in Sec.
+4.
+During the local training on client c, given the model parameter $\theta$, we aim to show that optimizing $\min_{\theta}L_{\text{task}}(z^{c*}_{I},z^{c*}_{T})$ can be reformulated as optimizing $\min_{\theta}\left(L_{\text{task}}(z_{I}^{c},z_{T}^{c})+\alpha\left(\|z_{I}^{c}-z^{c*}_{I}\|_{2}^{2}+\|z_{T}^{c}-z^{c*}_{T}\|_{2}^{2}\right)\right)$.
+Since $L_{\text{task}}(x,y)$ is $L$-lipschitz, we have:
+$\|L_{\text{task}}(x_{1},y_{1})-L_{\text{task}}(x_{2},y_{2})\|\leq L(\|x_{1}-x_{2}\|+\|y_{1}-y_{2}\|).$ (7)
+Hence, we have:
+$L_{\text{task}}(z^{c*}_{I},z^{c*}_{T})$ $\leq\|L_{\text{task}}(z^{c*}_{I},z^{c*}_{T})-L_{\text{task}}(z^{c}_{I},z^{c*}_{T})\|+\|L_{\text{task}}(z^{c}_{I},z^{c*}_{T})\|$
+$\leq L(\|z^{c}_{I}-z^{c*}_{T}\|+\|z^{c*}_{T}-z^{c*}_{T}\|)+\|L_{\text{task}}(z^{c}_{I},z^{c*}_{T})\|$
+$=L\|z^{c}_{I}-z^{c*}_{I}\|+\|L_{\text{task}}(z^{c}_{I},z^{c*}_{T})\|$
+$\leq L(\|z^{c}_{I}-z^{c*}_{I}\|+\|z^{c}_{T}-z^{c*}_{T}\|)+\|L_{\text{task}}(z^{c}_{I},z^{c}_{T})\|$
+$=L(\|z^{c}_{I}-z^{c*}_{I}\|+\|z^{c}_{T}-z^{c*}_{T}\|)+L_{\text{task}}(z^{c}_{I},z^{c}_{T})$
+Therefore, $L(\|z^{c}_{I}-z^{c*}_{I}\|+\|z^{c}_{T}-z^{c*}_{T}\|)+\|L_{\text{task}}(z^{c}_{I},z^{c}_{T})\|$ can be an upper bound of the original learning object $L_{\text{task}}(z^{c*}_{I},z^{c*}_{T})$.
+During local training, we add the L-2 regularization to guide the update of the local model.
+We also consider the contrastive loss to measure the distance between $(z^{c}_{T},z^{c}_{I})$ and $(z^{c*}_{T},z^{c*}_{I})$ in the ablation study.
+However, L-2 regularization brings better performance and matches our assumption.
+G.2 Discussion on Cross-Modal Alignment
+Here we illustrate why we model the cross-modal alignment as an alignment module that projects features of vision and language modalities to a shared representation space, and intends to minimize the distance between representations of the two modalities.
+Following the notation in F,suppose we have features $I,T$ of vision and language modalities, and the corresponding alignment module $g_{I}(\cdot),g_{T}(\cdot)$.
+Inspired by Huang et al.
+(2021b), suppose $g_{I}(\cdot),g_{T}(\cdot)$ are $L$-Lipschitz, we formulate the alignment degree of two domains as:
+$R_{\epsilon}=\mathbb{P}\left[u:\sup_{I=H_{I}(u),T=H_{T}(u)}\|g_{I}(I)-g_{T}(T)\|\geq\epsilon\right]$ (8)
+For a simple case following Huang et al. (2021b), the goal of the alignment module can be written as:
+$L_{\text{align}}(g_{I},g_{T})=\mathbb{E}_{x}\mathbb{E}_{I=H_{I}(u),T=H_{T}(u)}\|g_{I}(I)-g_{T}(T)\|^{2}$ (9)
+Using the in Theorem 2 in Huang et al. (2021b) by replacing the encoder to our $g(\cdot)$, we have the following bound:
+$R^{2}_{\epsilon}\leq\eta(\epsilon)^{2}\cdot\mathbb{E}_{u}\mathbb{E}_{I=H_{I}(u),T=H_{T}(u)\in A(x)}\|g_{I}(I)-g_{T}(T)\|^{2}=\eta(\epsilon)^{2}\cdot L_{\text{align}}(g_{I},g_{T})$ (10)
+,where $\eta(\epsilon)$ is a infite function that corresponding to threshold $\epsilon$.
+Therefore, we model the cross-modal alignment as the alignment module that minimizes the distance between the representations of two modalities.
+Appendix H Limitation
+The design of the proposed FedAID approach has taken it into serious consideration to mitigate the potential negative societal impact..
+For example, to tackle the data privacy issue, FedAID has avoided the transmission of any data or latent representations during all server-client communications, while only transferring model parameters and loss, thereby minimizing the risk of privacy leakage in federated learning as much as possible.
+However, our model may still be subject to privacy-related issues due to the communication of model parameters, which is a common problem faced by other federated learning methods as well.
+There are also research studies in this line to avoid privacy leakage from model parameters.
+For fairness, our method aims at optimizing the performance of the worst-case client to promote model equity across different cohorts, whichhas shown great generalizable ability on unseen data.
+This approach ensures that the pre-trained model not only performs well and fairly on the data from hospitals participating in the federated training but also demonstrates good performance on hospitals that did not participate in the training.
+Meanwhile we acknowledge that there may still exist limitations.
+Firstly, If we can resolve the trade-off between privacy protection and data representation communication, it may allow us to design more advanced algorithms to further enhance the performance of federated multimodal pre-training models.
+Furthermore, in this work we mainly consider the multimodal medical datasets of X-Ray imageeport due to the limitation of the public datasets.
+However, some rare diseases and more complicated modalities in various medical applications lack paired data, which] may bring other challenges for federated multimodal pretraining.